### v1

# Positional encoding is not the same as context: A study on positional encoding for Sequential recommendation

Abstract
The expansion of streaming media and e-commerce has led to a boom in recommendation systems, including Sequential recommendation systems, which consider the user’s previous interactions with items. In recent years, research has focused on architectural improvements such as transformer blocks and feature extraction that can augment model information. Among these features are context and attributes. Of particular importance is the temporal footprint, which is often considered part of the context and seen in previous publications as interchangeable with positional information. Other publications use positional encodings with little attention to them. In this paper, we analyse positional encodings, showing that they provide relative information between items that are not inferable from the temporal footprint. Furthermore, we evaluate different encodings and how they affect metrics and stability using Amazon datasets. We added some new encodings to help with these problems along the way. We found that we can reach new state-of-the-art results by finding the correct positional encoding, but more importantly, certain encodings stabilise the training.
1 Introduction
Sequential recommendation systems (SRS) have been booming within referral systems in the last few years since most of these systems are based on previous user behaviour and are present in almost all e-commerce, social media, and streaming platforms. Their effectiveness has led to various definitions of sequential recommendations and to their division into branches.
In general, sequential recommendation systems Fang et al. (2020), refer to recommendation systems (RS) where the interaction between the user and items is recorded sequentially. SRS are also related to session-based or session-aware recommendations. The latter two terms can be viewed as the sub-types of sequential recommendation when the interactions log is restricted to those made in the current session.
In this work, we adhere to the notation used in Fang et al. (2020) and use the term object to summarise the interactions. "Behavior object" means what the user interacts with, identified by an ID or a group of items. It may include text, images, and time. We will use "item(s)" for convenience. The interaction between items and users can use many different types of features. A typical split of those is attributes and context, where the attributes are features provided by the item, like a text description or an image, and the context refers to those features related to the interaction itself, like the time that a user tends to buy or search some streaming content.
A division based on the target task exists between the next-item recommendation and the next-basket recommendation Fang et al. (2020). In the next-item recommendation, a behaviour contains only one object (i.e., item), which could be a product, a song, a movie, or a location. In contrast, a behavior contains more than one object in next-basket recommendation.
When dealing with sequential systems, it is essential to consider the sequential nature of the data. Therefore, initial deep learning techniques utilised Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN) to address this issue. However, attentional models have become increasingly popular due to their effectiveness, leading to multiple approaches utilising this technique Wang et al. (2021). Currently, most state-of-the-art models use attentional networks.
To summarise, our contribution is fourfold:
  1. 1.
    We analyse existing work and highlight that more than the standard usage of $5$ runs is needed to ensure the reliability of the results.
  2. 2.
    We propose new encodings. Based on our first experiments with existing encodings, we propose new encodings to improve metrics and stability.
  3. 3.
    We perform a detailed analysis of the types of encodings, their features (in-head, affected heads, etc.) and their ablations to see which factors improve the stability and the metrics.
  4. 4.
    We found correlations between the sparsity of the dataset and the encodings to get the best results. This precise data dependence, together with our study, provides a road map for the usage of encodings for SRS with transformer architectures.
2 Related Work
From an architectural perspective, SRS has evolved over the last two decades. Traditional methods in the field applied classic ML techniques, such as KNN for item-based systems (Davidson et al. (2010), Linden et al. (2003)), or applications of Markov Chains for different orders (He et al. (2017), He et al. (2016b), He and McAuley (2016a)). A common approach due to the sparsity of the interactions in RS is those based on matrix factorisation like DeepFM Guo et al. (2017) or CFM Xin et al. (2019). Among the different DL architectures, RNNs can capture sequence information and represent items based on their position in the list. One of the earliest and best known was Cho et al. (2014). The first approach from DL came using RNN, the most common architecture then. We can find models such as GRU4Rec Hidasi et al. (2016) or DREAM Yu et al. (2016). Another of the first approaches via DL was based on CNN models that can capture local features and consider time information in the input layer (Tang and Wang (2018), Tuan and Phuong (2017)). Among these were Caesar Li et al. (2021) and NextItNet Yuan et al. (2018). One of the first approaches using attention models was NARM Li et al. (2017), an encoder-decoder framework that uses attention at the bottleneck. The introduction of transformers and self-attention in sequential models came from AttRec Zhang et al. (2018).
Over time, different types of information have been identified to predict what a user might buy next. There are three main types of models for this: attribute-aware Rashed et al. (2022), context-aware Zhou et al. (2020), and time-aware Li et al. (2020). Attribute-aware models focus on the information provided by the items themselves. Context-aware models take into account the characteristics of the user’s interaction. Time-aware models are a subgroup of context-aware models that use the timing of the user’s interactions.
Using transformers in architecture has led to the introduction of positional embeddings. By default, the attention heads work like bags of tokens, and the positional information needs to be added to understand the sequential order of the tokens. Some approaches use time-aware information to replace the position, like in the case of CARCA Rashed et al. (2022). While this approach achieved excellent results, it faces some limitations. Replacing the positional embeddings with time stamps assumes that the order of the items will be implicitly deduced from this stamp. This can be more challenging to the model than simply an explicit representation. Furthermore, recent research in NLP Chen et al. (2021) has shown that introducing positional information within the attention, instead of the first head input, produces better results, which can not be done through the time stamps.
One of the first models considering the sequential information was GRU4Rec Hidasi et al. (2016), which concentrates on extracting information from the sequence of items instead of the user being an anonymous user model. A successful model from the context-aware point of view was DeepFM Guo et al. (2017), which extracted through the combination of factorisation models and deep learning extracts information from the context. Like GRU4Rec, using a self-attention layer to balance short-term intent and long-term preference can be found in SASRec Kang and McAuley (2018a). The extraction through a sequence of transformer encoders produces an output, which is compared with the target items through dot-product. This dot-product approach combines information from past behaviour and target representation, which was improved by an encoder-decoder architecture, where the encoder extracts past information, and the decoder combines the embedding representation for the target. This produced one of the SOTA models, CARCA Rashed et al. (2022). Another path of improvement from SASRec came from training the bidirectional transformer to infer the item-item relationship from the user’s historical interactions using the Cloze task in BERT4Rec Sun et al. (2019). Another one was TiSASRec Li et al. (2020), which further improves SASRec by considering time intervals between items.
We will focus on the position encoding applied to the most common datasets among these papers. For a deeper review of the different encoding techniques, one can see Zhao et al. (2024). Since there is no systematic study, each work addresses the encoding differently. In AttRec Zhang et al. (2018), they "furnish the query and key with time information by positional embeddings" using absolute embeddings. SASRec used absolute learnable position embeddings in the input sequence. They also tried the fixed position embedding as used in Vaswani et al. (2017) but found that this led to worse performance. In CARCA, which will be our base model, they tried using absolute encodings again, giving even worse results than no positional encoding. This disparity of results merits a more detailed analysis as in the present work.
3 Positional encodings
Encoding Learnable Type Range Layers Concatenation Integration
Abs No APE Fix. Vector Encoding No Add
Abs + Con Yes APE Fix. Vector Encoding Yes Joint
Rotatory No APE Unl. Vector Encoding No Multi.
Rotatory + Con. Yes APE Unl. Vector Encoding Yes Joint
RMHA-4 No RPE 4 All No Multi.
Rope No APE Unl. All No Multi.
RopeOne No APE Unl. First No Multi.
Learnt Yes APE Fix. Vector Encoding No Add
Learnt + Con Yes APE Fix. Vector Encoding Yes Joint
None No N/A N/A 0 N/A N/A
Table 1: Encoding types: ’Type’ refers to absolute and relative. ’Learnable’ indicates whether the encoding parameters are learned during training. We also consider learnable to have a concatenation since there are learnable parameters. ’Range’ specifies the positions the encoding covers: ’Fix.’ fixed by the given length, ’Unl.’ means unlimited, while $4$ refers to the relative range. ’Layers’ represents the number of layers the encoding uses: "Vector Encoding" means they are added at the start by vectors (left in Fig: 1), All means in all the transformer blocks and First, at the first encoding block. ’Concatenation’ indicates whether there is a concatenation. ’Integration’ describes how the encoding is integrated with the model architecture: ’Add’ indicates a vector sum, ’Joint’ the concatenation, and ’Multi.’ for multiplication of matrices for those in the heads or element-wise for the others. Finally, None refers to no encoding.
Figure 1: APE vs RPE : While APE encodings introduce the positional information as vectors before the $V$, $Q$ and $K$, projections, RPE add this information at the coefficient level in $K$ and $Q$. Parts affected by the position information appear in red.
3.1 Positional encodings
This section introduces the different types of encoding techniques used in this work. For a deeper review of the various encoding methods, check Zhao et al. (2024). Following Zhao et al. (2024), there are two prominent families of encoding methods. First, we have absolute positional encoding (APE ). These encodings are added before the first Transformer block, and the positional information provided is absolute, from the initial token to the last one. Second, there are the relative encoding methods (RPE ), which provide information about those items around the target item. These encodings are implemented inside the transformer’s head, more precisely in the Queries and Keys. In contrast to the common usage of encoding in Transformers for NLP tasks in RS, the encoding is reversed; the first item in the encoding order is the last one, and padding is applied on the left side Kang and McAuley (2018a). We chose this approach because the most relevant items are usually the ones that were recently visited or clicked. Due to the sinusoidal functions used in APE , the most pertinent items might be at the start because the assigned frequency becomes less distinguishable the later an item appears. We did some preliminary experiments to check if this is also right for another encoding, and the reversed encoding always produced better results, so in what follows, all the encoding will be considered as reversed.
Self-attention
enables models to weigh different parts of input sequences differently, allowing them to focus on relevant information during processing. The self-attention mechanism was introduced in Vaswani et al. (2017), which laid the foundation for the transformer architecture.
In self-attention, the input sequence is typically represented as a set of vectors, where each vector corresponds to a position in the sequence. This allows each position to attend to all other positions and captures dependencies and relationships within the sequence. For a given dimensionality $d$ of the item representation and a hidden dimension $H<d$, the attention weight assigned to each position is computed based on the similarity (dot product) between the query $Q\in\mathbb{R}^{d\times H}$, key $K\in\mathbb{R}^{d\times H}$, and value $V\in\mathbb{R}^{d\times H}$ matrices associated with the positions. These learnable matrices are derived from linear transformations of the original input $X\in\mathbb{R}^{n\times d}$ by multiplication, where $n$ is the number of items in the sequence.
The attention weight ($\alpha_{ij}$) assigned to position $i$ for attending to position $j$ is computed using the following formula:
$\alpha_{ij}=\frac{{\exp((X_{i}\,Q)\cdot(X_{j}\,K)^{T}/\sqrt{d})}}{{\sum_{t=1}^{n}\exp((X_{i}\,Q)\cdot(X_{t}\,K)^{T}/\sqrt{d})}}$ (1)
Here, $Q_{i}$ and $K_{j}$ represent the query and key matrices associated with positions $i$ and $j$, respectively. The dot product is scaled by $\sqrt{d}$ to prevent vanishing gradient problems. The resulting attention weights are then used to compute a weighted sum of the value matrices ($V_{j}$) associated with all positions:
$\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}X\cdot V_{j}$ (2)
This weighted sum, representing the attended information, is then used as the output of the self-attention mechanism. The process is repeated for all positions in the sequence.
Absolute Positional Encoding (APE )
Absolute positional encoding adds fixed values to the embeddings of items based on their absolute position in the sequence. This means all the items have the same initial token as a reference. The most common approach is based on trigonometry, like in BERT Devlin et al. (2019), the so-called sinusoidal encoding. Absolute encoding is added at the input layer as part of the item encoding, 1. The most common practice is to add a vector encoding the position.
The absolute positional encoding for a token at position $i$ in a sequence of length $N$ can be calculated using trigonometric functions such as sine and cosine. A common formula is:
$PE_{(pos,2i)}=\sin\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$ (3)
$PE_{(pos,2i+1)}=\cos\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$ (4)
where $pos$ is the position of the item, $i$ is the dimension, and $d$ is the dimension of the embedding. This encoding is usually added to the original embedding, and we will refer to it along the text as absolute encoding (AE). For a given item embedding $x$ at position $pos$, we replace it as follows:
$x=x+PE_{(pos,*)}$ (5)
Another way of making the positional information more explicit and separated from the rest of the context and attributes is by concatenation, where we divide the dimension of the embedding into the context and attribute information and the positional encoding information:
$x=Concat\,(x,PE_{(pos,*)})$ (6)
We apply a Linear layer right after to recover the original dimension and do not change the number of parameters in the model. We will refer to them as concat absolute encoding (CAE). We added the position information orthogonally to make the model gradually rely on the position, producing more stable results, Section 5.
Learnable Encoding
Learnable position embeddings take a different approach by treating the positional information as parameters that can be updated during training. Instead of relying on predefined patterns, the model learns the optimal representation of positions based on the data context. We will denote these embeddings as learnable encoding (LE). In this case, we add a positional encoding, which is merely a learnable vector. We also apply a concatenation approach denoting as concatenated learnable encoding (CLE). In this case, their results will be generally worse than the classic residual connection.
Relative Positional Encoding (RPE )
A disadvantage of using APE is that it requires a fixed length of the input sequence and does not directly capture the relative positions of one item to another, whether the other is further ahead or backward. Relative position encoding (RPE ) was proposed to solve these issues, 1. Shaw et al. (2018) uses relative position encoding (RPE ) instead of APE and adds the position embeddings to the key and optional value projections instead of the input. A simplified version deleting the position embeddings in value projections in TranformersXL Dai et al. (2019) showed better performance on language modelling tasks. The length of the user’s session or action log varies significantly from user to user, suggesting that these encoding techniques may be of value to SRS.
The relative positional encoding can be computed using trainable parameters. The specific formula may depend on the implementation; in our case, we follow the one in Shaw et al. (2018), where the position is given inside the attention. The relation between two inputs $X_{i}$ and $X_{j}$ is given by $a_{ij}^{V},a_{ij}^{K}\in\mathbb{R}^{d_{a}}$, for values and keys. This information can be shared among heads, but we consider the case in which information is not shared.
For the relative case, the attention weight ($\alpha_{ij}$) can be written as:
$\frac{{\exp((X_{i}\,Q)\cdot(X_{j}\,K+a_{ij}^{K})^{T}/\sqrt{d})}}{{\sum_{t=1}^{n}\exp((X_{i}\,Q)\cdot(X_{j}\,K+a_{ij}^{K})^{T}/\sqrt{d})}}$ (7)
We also introduce the position at the value product as follows:
$\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}(X\cdot V_{j}+a_{ij}^{V})$ (8)
Relative positional encoding can be especially useful when dealing with long sequences, as it allows the model to generalise better to different distances between items and in cases where the local position information is more critical. As a drawback, since it has to be applied to each attention head, it tends to be slower Shaw et al. (2018). We will denote these embeddings as RMHA-4 , where we bound the relative indexes to a distance of $4$. We do not apply a concatenated version in this case.
Rotary Encoding (Rope and Rotatory )
Unlike absolute or relative positional encoding, which focuses on representing positions in a fixed coordinate system or relative to other items in a sequence, rotary encoding leverages rotational transformations to encode position information. "Rotary Positional Encoding" typically refers to a specific technique used in positional encoding, Su et al. (2024). We will use the notation Rope to refer to them.
The motivation behind Rope is that the orientation or rotation of positional embeddings can provide valuable information about the relationship between items in a sequence. This approach is particularly relevant in scenarios where the inherent order or arrangement of items is not straightforward and can benefit from a more flexible representation, like a user searching for more than one item in the same session. These embeddings capture both linear order and rotational relationships within the sequence. Rotary position encodings apply a rotation to the Query and Key matrices, which preserves the dot product among the vectors.
In the two dimensional case for a given input vector $x$ the encoding function $f$ for given angle $\theta$ is defined as follows:
$f_{\theta}(\mathbf{x})=\mathbf{R}(\theta)\cdot\mathbf{W}\cdot\mathbf{x}$ (9)
Here, $\mathbf{R}(\theta)$ represents the rotation matrix parameterized by an angle $\theta$, expressed using sines and cosines, and $\mathbf{W}$ is either $K$ or $Q$. More explicit, for the $K$ case we have:
$\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\cdot\begin{bmatrix}K_{11}&K_{12}\\ K_{21}&K_{22}\end{bmatrix}\cdot\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$ (10)
The straightforward generalisation of this to higher dimensionality divides the vector dimension into two-dimensional chunks and applies the rotation to each of them. Since this implementation is costly, the actual implementation expresses this rotation as two vector multiplications and one vector addition, as RoFormer Su et al. (2024).
Since we want to analyse the effect of the rotational position for sinusoidal, we propose here a new encoding, Rotatory , which is concatenated version Rotatory + Con. . Instead of directly applying sinusoidal functions, rotary encodings introduce an additional rotation mechanism by separating the angles into sine and cosine components and then alternating them across the embedding dimension. By alternating sine and cosine embeddings, the model can capture different phases of positional information, effectively introducing a rotational component to the encoding.
This rotation helps mitigate potential issues, such as vanishing gradients that can arise with traditional positional encoding, which is especially important for stability. So, while the term "rotary" might not directly imply physical rotation, it refers to the rotational mechanism introduced in the encoding process to enhance the representation of positional information in Transformer models. The implementation is similar to that from Rope , but applying the rotation matrices to the vector encodings added to the input items.
Dataset Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg
Beauty silu Rotatory-Longer 0.0001 61.87 1.04 42.60 0.98
Beauty silu Rotatory 0.0001 61.72 1.14 42.54 1.52
Beauty silu Rotatory + Con. 0.0001 61.68 1.04 42.31 0.97
Beauty leaky Rotatory + Con. 0.0001 61.16 0.77 42.83 0.48
Men silu RMHA-4-Longer 0.0001 70.13 0.42 46.41 4.18
Men silu RMHA-4 0.0001 69.70 0.64 43.75 1.44
Men leaky None 0.1000 69.69 1.28 57.94 1.69
Men leaky RMHA-4-Longer 0.0001 68.72 1.47 43.46 0.79
Fashion silu RMHA-4 0.0001 77.26 0.24 49.75 2.33
Fashion silu RMHA-4-Longer 0.0001 77.00 0.48 49.40 0.54
Fashion leaky RMHA-4 0.0001 76.46 0.25 49.49 2.18
Fashion leaky RMHA-4-Longer 0.0001 76.20 0.51 49.85 3.92
Games leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79
Games silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55
Games leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93
Games silu Learnt NaN 79.82 1.88 54.30 2.95
Table 2: The table displays the top $4$ results for each dataset on cases with a standard deviation for $HIT$ below $3$.
Dataset Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg
Beauty leaky Abs + Con 0.0001 67.93 4.17 48.71 3.59
Beauty silu Learnt 0.0001 67.62 6.95 46.33 7.08
Beauty silu Abs 0.0001 67.27 10.59 45.34 10.75
Beauty silu Abs + Con 0.0001 65.87 5.23 45.61 5.12
Men leaky Learnt + Con 0.1000 73.86 7.18 58.89 8.15
Men leaky Learnt 0.0001 72.08 4.44 56.55 9.98
Men leaky Learnt + Con 0.0001 71.34 7.13 57.92 13.07
Men leaky Rotatory + Con. 0.1000 70.92 3.92 59.84 4.88
Fashion silu RMHA-4 0.0001 77.26 0.24 49.75 2.33
Fashion silu RMHA-4-Longer 0.0001 77.00 0.48 49.40 0.54
Fashion leaky RMHA-4 0.0001 76.46 0.25 49.49 2.18
Fashion leaky RMHA-4-Longer 0.0001 76.20 0.51 49.85 3.92
Games leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79
Games silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55
Games leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93
Games silu Learnt NaN 79.82 1.88 54.30 2.95
Table 3: he table displays the top $4$ results for each dataset on cases with a standard deviation for HIT below $12$.
4 Experimental Setup
4.1 Datasets
We selected the Amazon dataset to assess the proficiency of various encoding types. For this purpose, we utilised four different real-world datasets extracted from product reviews on Amazon.com. These datasets have been diverse and widely used for SRS under the leave-one-out protocol (He and McAuley (2016b), Kang and McAuley (2018b), Hou et al. (2019), Wu et al. (2020), Steck (2019) and Zhou et al. (2020)). Since multiple versions of these datasets exist, we opted for the preprocessed versions from Rashed et al. (2022).
These datasets encompass various product categories, providing a diverse and comprehensive source for training and evaluating recommendation models. They provide essential features for each interaction, such as user IDs, item IDs, timestamps, and additional vector contextual information. As we can see in Tables 5, and 4, they provide diversity in the sparsity, which is going to be a key point in our analysis, but also on the discreteness of the features.
Dataset Users Items Interactions Attributes
Men 34,244 110,636 254,870 2,048
Fashion 45,184 166,270 358,003 2,048
Games 31,013 23,715 287,107 506
Beauty 52,204 57,289 394,908 6,507
Table 4: Dataset Statistics.
1.Beauty
(Wu et al. (2020), Zhou et al. (2020) and Kang and McAuley (2018b)) The dataset includes discrete and categorical attributes for all beauty products, including fine-grained categories and brands. Mainly categorical and discrete features.
2.Video Games
(Wu et al. (2020) and Kang and McAuley (2018b)) The Video Games sub-dataset includes user interactions, reviews, and product details specific to the video game category as price, brand and categorical features. Most of the attributes here are discrete and categorical.
3.Men
(He and McAuley (2016b)) The men’s dataset encompasses a comprehensive collection of items falling under men’s clothing. The attributes are dense vectors from image-based features extracted from the last layer of a ResNet50 Deng et al. (2009) on the ImageNet dataset He et al. (2016a).
3.Fashion
(He and McAuley (2016b) and Hou et al. (2019)) It contains 6 categories for men’s and women’s clothing. The features were extracted as dense using the same ResNet50 approach.
4.2 Models and Training
To implement CARCA, we converted TensorFlow code from the original model into Pytorch [^1] https://github.com/researcher1741/Position_encoding_SRS. We extracted the implementation in Papso (2023) for the part of the loss. The original version in TensorFlow was written on top of previous models written in TensorFlow. We found that these implementations check the metrics every $20$ epoch, assuming that the overfitting starts later. This is probably because some models need more than $1000$ epochs, and the cost of measurement becomes significant. We measure the metrics in a logarithmic manner, giving rise to slightly better results in some cases. Our implementation can be found here .
As in previous works (Rashed et al. (2022), Kang and McAuley (2018a), etc.), we use an ADAM optimiser to minimise the binary cross-entropy loss of the CARCA model while masking the padded items to prevent them from contributing to the loss function. For a given sequence of items interactions for a user $u$ as $\{i^{u}_{1},\ldots,i^{u}_{N}\}$, we create an input list as $I^{u+}=\{i^{u}_{1},\ldots,i^{u}_{N-1}\}$ by deleting the last item of the list, a positive target list as $T^{u+}=\{i^{u}_{2},\ldots,i^{u}_{N}\}$ by shifting one the input list, and a negative target list as $T^{u-}=\{i^{rand_{1}},\ldots,i^{rand_{N}-1}\}$ generated by random negative items. Then the loss is given by:
$-\sum_{u\in U}\sum_{r\in T^{u+}\cup T^{u-}}\left[Y_{r}\log(\hat{Y}_{r})+(1-Y_{r})\log(1-\hat{Y}_{r})\right]$
Here, $U$ represents the set of users, $T^{u+}\cup T^{u-}$ denotes the union of positive and negative items for user $u$, $Y_{r}$ is the observed interaction label for item $r$, and $\hat{Y}_{r}$ is the predicted probability of interaction with item $r$.
4.3 Metrics
We use two commonly used metrics for Next-Item Recommendations/SRS. Hit Rate measures the proportion of recommended items that overlap with the user’s actual interactions. Hit@10 counts the fraction of times the ground truth next item is among the top 10 items.
The second metric is Normalized Discounted Cumulative Gain (NDCG). NDCG considers the relevance of items and their positions in the recommended list.
$NDCG=\frac{DCG}{IDCG}$ (11)
Where $DCG$ (Discounted Cumulative Gain) is calculated as the sum of the relevance scores of items at each position in the ranked list, discounted by their position, and $IDCG$ (Ideal Discounted Cumulative Gain) represents the maximum possible $DCG$ for a perfect ranking.
$DCG=\sum_{i=1}^{n}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$ (12)
$IDCG=\sum_{i=1}^{n}\frac{2^{max(rel_{i},0)}-1}{\log_{2}(i+1)}$ (13)
Here, $rel_{i}$ is the item’s relevance at position $i$ in the ranking.
5 Results
We present all the results for all the datasets and encodings in Appendix A.1. Due to space limitations, sub-tables were extracted from them.
Together with position encoding, we wanted to analyse other factors that could affect stability. In CARCA Rashed et al. (2022), they found $0.0001$ as the best upper bound ($nmax$ in our Tables) for the encoding vectors in all the datasets except for Games. We try different maximum values for all the encoding vectors ($0.1$ and $0.0001$) to see if their choice depends on the encoding. Except for some exceptions, the upper bound $0.0001$ produces better results. We also tried the activation function $silu$ instead of the $leakyrelu$. The results are slightly better regarding stability but without a strong correlation.
5.1 Stability
A common practice among the papers is to present an average of $5$. Our first experiments showed a significant difference between each one of the runs. We decided to analyse the deviation for the different trainings. The deviation and mean can be found in all the training tables. We also decided to compute the $95\%$ significance confidence interval and present it with the confidence interval in the complete tables results in Appendix A.1.
Inter-Dataset Deviation:
The stability varies from one dataset to another, being the Beauty dataset the most stable with an average Hit deviation of $0.9$ and NDCG deviation of $0.87$. Meanwhile, the maximum deviations are obtained by fashion. We can see that those datasets with a lower sparsity tend to be more stable, as do the Beauty and Games datasets, while those with a higher sparsity are less stable, as are the Men and Fashion datasets.
Inter-Encoding Deviation:
In Table 6 and Table 12 in the Appendix, we see that some encodings produce more stable results than others. More precisely, the concatenated version for Abs , Learnt and Rotatory + Con. generally present more stable results. Furthermore, for the more stable case of $0.0001$ upper bound, in Table 6, Abs and Learnt are still significantly stable. At the same time, our Rotatory encoding produces similar results to the concatenated version. This could be because the average deviation is similar to no encoding, None . The most notable case is RMHA-4 , which, combined with the upper bound $0.0001$, gives the most stable results, around $0.51$, compared with the no-encoding case, $2.36$. This pattern is preserved to the general upper bound, suggesting that this encoding provides excellent stability. As we will see below, this encoding is worth trying in cases of high sparsity.
Encoding Ablation:
We can appreciate that Learnt and the sinusoidal Abs tend to have a similar deviation, higher than not using any encoding. To know if the stability is provided by the fact that RMHA-4 introduce positional information in each transformer or being a RPE head, we trained Rope . The results show that it is a mid between RMHA-4 on the one hand and Learnt and Abs on the other hand. It does not add or reduce stability for None . We wanted to know if being Rope in a middle point is due to being in all the transformer blocks while not being a APE encoding. To this goal, we added RopeOne , an ablation which applies Rope but only in the first block. The results are not far away from each other. We can see in Table 12 that RopeOne is even slightly better, while in Table 6, they are almost the same, in this case, slightly worse. Then, we decided to create a new encoding, Rotatory , which uses the rotatory properties but is not in the attention heads but a vector encoding. The results show the same stability as Rope and its ablation.
The stability of the RMHA-4 encoding can also be appreciated in the losses. In Appendix A.5.
Dataset Hit Dev. NDCG Dev. CI-length Density/Sparsity-1
Men 3.9143 8.2742 5.6686 $56.727\times 10^{-}$ 5
Fashion 4.87 9.7814 9.2257 $54.765\times 10^{-}$ 5
Games 1.124 0.952 1.616 $43.904\times 10^{-}$ 4
Beauty 0.935 0.87 1.22 $41.320\times 10^{-}$ 4
Table 5: Deviations among the different datasets without Encoding. This pattern is generally preserved along the different encodings.
encoding dev HIT dev NDCG runs CI-length
Abs 5.43 8.25 9.38 6.71
Abs + Con 3.52 5.56 7.12 5.35
Learnt 4.79 7.64 7.88 7.05
Learnt + Con 3.17 6.36 6.62 4.42
None 2.35 5.26 8.00 3.29
RMHA-4 0.51 1.08 6.00 0.88
Rope 2.38 5.45 5.00 4.04
Rotatory 2.46 5.01 8.38 3.48
Rotatory + Con. 2.48 5.18 6.75 3.49
RopeOne 2.82 6.54 4.29 5.46
Rope-Longer 3.33 8.15 5.25 5.99
Rotatory-Longer 1.05 1.04 4.67 2.08
Rotatory + Con.-Longer 0.99 1.24 6.00 1.59
RMHA-4-Longer 0.71 4.00 6.00 1.14
Table 6: Deviations among the different encodings. Relative encoding in all the heads, RMHA-4 , presents more stability.
5.2 Best Results
Which is the best encoding?
It depends on the spacing and the definition that one manages to get a good result. If we look at the table 7, we see that this dataset has excellent stability, but in cases like Learnt with $0.1$, with leakyrelu, the model gets over $74$ in HIT, but the $95\%$ confidence interval is ($60.74$, $88.24$). In other words, applying this approach cements the results of what was studied here Picard (2021); in some cases, it is better to get a good set of seeds than a suitable method.
Maximum deviation of $12$:
If we consider a maximum of $12$ as deviation, we get Table 3. In CARCA, they did an ablation study on the Men Dataset, showing that using Abs moves towards worse results. They produced similar results for Men, but both had a deviation of around $6$, so the results from the ablation were more a matter of initialisation. In this case, Men and Beauty get their best results with Learnt -like and Abs -like encodings, respectively. But they are unstable. In contrast, Fashion and Games get their top values with RMHA-4 and Rotatory , respectively, which are stable.
Maximum deviation of $3$:
If we reduce the maximum deviation to $3$, we get some results that we consider the right results. As we can see in Table 2, the best encodings are RMHA-4 and Rotatory (or its concatenated version), and they correlate with the sparsity. Those with a high sparsity benefit more from the extra stability provided by RMHA-4 , while those with a lower sparsity, Beauty and Games, get their best through a purely rotatory approach Rotatory . We can see that even considering only those which are stable, the results obtained in all the datasets are new SOTA.
Encoding "Ablation":
We created some ablations from existing encoding techniques to analyse where the stability came from. So far, our experiments used almost the same hyperparameters as the original CARCA. Looking again at the losses, we found that some of the top results did not converge, i.e., the loss and the metrics were improving. Thus, we decided to extend some trainings with more epochs to see if this produced better results. We added $400$ epochs to the default values (11). The improvements were not significant in this regard.
6 Conclusion
Our main conclusion is that encodings should not be considered minor in this model family. On the contrary, they are a powerful tool, and depending on the data set, a couple of possible better choices can be pre-selected. We have shown that the higher the sparsity, the better the results, as they stabilise the training until the best results are achieved. On the other hand, when the set density is sufficient, our rotatory encodings are a good choice or a learnable encoding. Also, at publication time, we have achieved SOTA results using only the encodings. And these results are, in most cases, more reliable than previous SOTA results.
7 Future work
During the encoding analysis, we found some correlation between the bound used for the encoding and the stability. Previous works have experimented with very low bounds, usually below $0.001$. We would like to go deep in this direction and analyse the behaviour of the metric used for encoding and its bounds.
8 Limitations
Despite the details of this study, it was carried out using a transformer architecture. Today, the latest models in the field are based on this architecture or GCNs. Using other architectures may affect this study and the conclusions on using encodings. The fact that the choice depends on the dataset can give a first idea of the possible encodings in case of using different architectures.
Appendix A Appendix
A.1 Experiments
In this section, we present all our results for the four datasets, thereby completing the missing experiments from the sub-tables in the main corpus. In tables 7, 8, and 9, we can find Rope-Longer , which is not presented in Table 1. These are just a few cases in which we extend the number of epochs since we found that for Rope , it takes longer to reach its maximum.
Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
leaky None 0.0001 56.00 0.96 37.93 0.65 9 (55.37, 56.63) 1.26
silu None 0.0001 57.03 0.83 38.42 0.68 9 (56.49, 57.57) 1.08
leaky None 0.1000 49.02 1.18 30.73 1.21 9 (48.25, 49.79) 1.54
silu None 0.1000 49.69 0.77 31.35 0.94 9 (49.19, 50.19) 1.00
leaky Learnt + Con 0.0001 56.36 0.65 37.94 1.54 3 (55.62, 57.1) 1.48
silu Learnt + Con 0.0001 56.69 0.51 37.82 0.65 3 (56.11, 57.27) 1.16
leaky Learnt + Con 0.1000 49.29 0.71 30.98 0.99 3 (48.49, 50.09) 1.60
silu Learnt + Con 0.1000 49.85 1.18 31.42 1.48 3 (48.51, 51.19) 2.68
leaky Abs + Con 0.0001 67.93 4.17 48.71 3.59 5 (64.27, 71.59) 7.32
silu Abs + Con 0.0001 65.87 5.23 45.61 5.12 5 (61.29, 70.45) 9.16
leaky Abs + Con 0.1000 51.42 1.41 32.10 0.46 3 (49.82, 53.02) 3.20
silu Abs + Con 0.1000 54.54 3.33 35.43 3.34 5 (51.62, 57.46) 5.84
leaky Rotatory + Con. 0.0001 61.16 0.77 42.83 0.48 3 (60.29, 62.03) 1.74
silu Rotatory + Con. 0.0001 61.68 1.04 42.31 0.97 3 (60.5, 62.86) 2.36
leaky Rotatory + Con. 0.1000 50.51 1.59 32.25 1.76 3 (48.71, 52.31) 3.60
silu Rotatory + Con. 0.1000 50.88 1.00 32.70 0.87 3 (49.75, 52.01) 2.26
leaky Learnt 0.0001 59.83 3.71 39.85 3.11 3 (55.63, 64.03) 8.40
silu Learnt 0.0001 67.62 6.95 46.33 7.08 5 (61.53, 73.71) 12.18
leaky Learnt 0.1000 74.49 17.19 51.36 22.65 6 (60.74, 88.24) 27.50
silu Learnt 0.1000 61.75 7.84 40.02 7.34 10 (56.89, 66.61) 9.72
silu Rotatory-Longer 0.0001 61.87 1.04 42.60 0.98 2 (60.43, 63.31) 2.88
silu Rotatory-Longer 0.1000 49.33 2.08 30.89 2.28 2 (46.45, 52.21) 5.76
leaky Abs 0.0001 63.65 8.36 43.68 8.36 13 (59.11, 68.19) 9.08
silu Abs 0.0001 67.27 10.59 45.34 10.75 15 (61.91, 72.63) 10.72
leaky Abs 0.1000 57.25 10.84 37.65 10.73 12 (51.12, 63.38) 12.26
silu Abs 0.1000 61.23 15.24 39.14 11.78 16 (53.76, 68.7) 14.94
leaky RMHA-4 0.0001 56.29 0.81 37.66 1.14 3 (55.37, 57.21) 1.84
silu RMHA-4 0.0001 57.00 0.35 37.22 0.32 3 (56.6, 57.4) 0.80
leaky RMHA-4 0.1000 51.30 0.32 33.38 0.15 3 (50.94, 51.66) 0.72
silu RMHA-4 0.1000 51.35 0.42 33.04 0.12 3 (50.87, 51.83) 0.96
leaky RopeOne 0.0001 55.33 1.22 37.89 1.24 3 (53.95, 56.71) 2.76
silu RopeOne 0.0001 57.25 1.16 38.76 1.01 3 (55.94, 58.56) 2.62
leaky RopeOne 0.0001 55.95 0.21 38.41 0.65 3 (55.71, 56.19) 0.48
silu RopeOne 0.0001 56.23 0.63 37.69 0.55 3 (55.52, 56.94) 1.42
leaky Rope 0.0001 56.30 0.92 38.60 0.82 3 (55.26, 57.34) 2.08
silu Rope 0.0001 56.59 0.72 37.21 0.98 3 (55.78, 57.4) 1.62
leaky Rope 0.1000 50.04 0.80 31.91 0.89 3 (49.13, 50.95) 1.82
silu Rope 0.1000 49.15 0.71 31.31 1.01 3 (48.35, 49.95) 1.60
leaky Rotatory 0.0001 58.69 1.31 40.83 0.71 5 (57.54, 59.84) 2.30
silu Rotatory 0.0001 61.72 1.14 42.54 1.52 5 (60.72, 62.72) 2.00
leaky Rotatory 0.1000 49.11 1.17 30.57 1.28 5 (48.08, 50.14) 2.06
silu Rotatory 0.1000 49.40 0.96 31.14 1.04 5 (48.56, 50.24) 1.68
Table 7: Beauty results with different positional encodings. Original results from Rashed et al. (2022) at the end.
Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
leaky None 0.0001 65.44 3.54 46.46 9.99 9 (63.13, 67.75) 4.62
silu None 0.0001 70.38 1.82 53.34 7.00 6 (68.92, 71.84) 2.92
leaky None 0.1000 54.04 7.99 38.82 9.53 3 (45.0, 63.08) 18.08
silu None 0.1000 61.53 11.84 48.13 14.97 3 (48.13, 74.93) 26.80
leaky Learnt + Con 0.0001 69.31 7.74 50.73 12.02 14 (65.26, 73.36) 8.10
silu Learnt + Con 0.0001 69.81 2.03 49.49 7.59 7 (68.31, 71.31) 3.00
leaky Learnt + Con 0.1000 65.29 7.61 52.91 9.70 10 (60.57, 70.01) 9.44
silu Learnt + Con 0.1000 66.15 7.47 50.59 9.46 14 (62.24, 70.06) 7.82
leaky Abs + Con 0.0001 67.30 4.48 52.26 9.94 10 (64.52, 70.08) 5.56
silu Abs + Con 0.0001 70.83 1.64 59.93 1.95 6 (69.52, 72.14) 2.62
leaky Abs + Con 0.1000 64.93 10.58 52.21 13.11 16 (59.75, 70.11) 10.36
silu Abs + Con 0.1000 65.14 8.20 51.71 11.06 13 (60.68, 69.6) 8.92
leaky Rotatory + Con. 0.0001 69.13 2.80 55.33 7.69 9 (67.3, 70.96) 3.66
silu Rotatory + Con. 0.0001 71.16 1.38 48.61 8.62 7 (70.14, 72.18) 2.04
leaky Rotatory + Con. 0.1000 68.61 2.76 57.08 3.72 7 (66.57, 70.65) 4.08
silu Rotatory + Con. 0.1000 65.23 7.76 52.99 9.31 11 (60.64, 69.82) 9.18
leaky Learnt 0.0001 68.37 5.75 51.41 12.43 13 (65.24, 71.5) 6.26
silu Learnt 0.0001 67.63 5.75 46.36 11.95 13 (64.5, 70.76) 6.26
leaky Learnt 0.1000 64.99 9.00 47.66 14.48 16 (60.58, 69.4) 8.82
silu Learnt 0.1000 63.94 9.49 46.44 10.58 16 (59.29, 68.59) 9.30
leaky RMHA-4-Longer 0.0001 76.20 0.51 49.85 3.92 6 (75.79, 76.61) 0.82
silu RMHA-4-Longer 0.0001 77.00 0.48 49.40 0.54 6 (76.62, 77.38) 0.76
leaky Rope-Longer 0.0001 68.63 3.97 51.05 11.71 5 (65.15, 72.11) 6.96
silu Rope-Longer 0.0001 71.50 1.22 49.69 6.41 7 (70.6, 72.4) 1.80
leaky Abs 0.0001 66.39 4.61 51.99 9.51 10 (63.53, 69.25) 5.72
silu Abs 0.0001 69.52 4.76 49.72 10.39 10 (66.57, 72.47) 5.90
leaky Abs 0.1000 59.75 9.74 44.73 12.86 15 (54.82, 64.68) 9.86
silu Abs 0.1000 58.39 8.48 41.86 11.31 15 (54.1, 62.68) 8.58
leaky RMHA-4 0.0001 76.46 0.25 49.49 2.18 6 (76.26, 76.66) 0.40
silu RMHA-4 0.0001 77.26 0.24 49.75 2.33 6 (77.07, 77.45) 0.38
leaky RMHA-4 0.1000 60.63 5.14 40.75 8.06 11 (57.59, 63.67) 6.08
silu RMHA-4 0.1000 63.61 4.41 39.48 7.97 11 (61.0, 66.22) 5.22
leaky RopeOne 0.0001 73.03 6.52 62.96 8.59 3 (65.65, 80.41) 14.76
silu RopeOne 0.0001 71.70 1.82 51.07 11.17 3 (69.64, 73.76) 4.12
leaky RopeOne 0.0001 69.27 5.02 54.24 11.80 6 (65.25, 73.29) 8.04
silu RopeOne 0.0001 70.49 1.66 49.17 10.70 6 (69.16, 71.82) 2.66
leaky Rope 0.0001 66.73 4.81 47.49 12.62 5 (62.51, 70.95) 8.44
silu Rope 0.0001 71.40 1.75 52.83 8.88 5 (69.87, 72.93) 3.06
leaky Rope 0.1000 61.24 8.78 48.13 10.86 6 (54.21, 68.27) 14.06
silu Rope 0.1000 62.79 8.79 49.27 10.69 8 (56.7, 68.88) 12.18
leaky Rotatory 0.0001 68.95 2.90 53.66 9.58 15 (67.48, 70.42) 2.94
silu Rotatory 0.0001 70.87 1.22 48.23 7.43 15 (70.25, 71.49) 1.24
leaky Rotatory 0.1000 62.02 10.97 48.79 13.69 18 (56.95, 67.09) 10.14
silu Rotatory 0.1000 62.68 8.53 49.41 10.86 15 (58.36, 67.0) 8.64
leaky None 0.0001 59.1 38.1 5
Table 8: Fashion Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
leaky None 0.0001 65.19 6.37 48.99 11.96 9 (61.03, 69.35) 8.32
silu None 0.0001 64.90 3.06 47.51 9.84 6 (62.45, 67.35) 4.90
leaky None 0.1000 69.69 1.28 57.94 1.69 3 (68.24, 71.14) 2.90
silu None 0.1000 65.66 0.89 53.12 0.67 3 (64.65, 66.67) 2.02
leaky Learnt + Con 0.0001 71.34 7.13 57.92 13.07 7 (66.06, 76.62) 10.56
silu Learnt + Con 0.0001 66.64 4.49 48.38 12.61 7 (63.31, 69.97) 6.66
leaky Learnt + Con 0.1000 73.86 7.18 58.89 8.15 7 (68.54, 79.18) 10.64
silu Learnt + Con 0.1000 62.18 7.83 47.82 10.83 7 (56.38, 67.98) 11.60
leaky Abs + Con 0.0001 70.84 5.13 57.63 9.89 9 (67.49, 74.19) 6.70
silu Abs + Con 0.0001 65.55 4.59 49.37 10.61 6 (61.88, 69.22) 7.34
leaky Abs + Con 0.1000 67.31 6.18 54.49 9.17 9 (63.27, 71.35) 8.08
silu Abs + Con 0.1000 63.33 7.50 48.62 10.54 9 (58.43, 68.23) 9.80
leaky Rotatory + Con. 0.0001 70.64 5.29 59.30 6.98 10 (67.36, 73.92) 6.56
silu Rotatory + Con. 0.0001 67.18 5.82 50.27 13.21 10 (63.57, 70.79) 7.22
leaky Rotatory + Con. 0.1000 70.92 3.92 59.84 4.88 6 (67.78, 74.06) 6.28
silu Rotatory + Con. 0.1000 63.88 7.42 50.69 9.75 10 (59.28, 68.48) 9.20
leaky Learnt 0.0001 72.08 4.44 56.55 9.98 7 (68.79, 75.37) 6.58
silu Learnt 0.0001 66.18 7.81 46.61 11.21 8 (60.77, 71.59) 10.82
leaky Learnt 0.1000 67.71 9.94 49.97 9.10 8 (60.82, 74.6) 13.78
silu Learnt 0.1000 67.15 12.19 51.53 14.44 8 (58.7, 75.6) 16.90
leaky RMHA-4-Longer 0.0001 68.72 1.47 43.46 0.79 3 (67.06, 70.38) 3.32
silu RMHA-4-Longer 0.0001 70.13 0.42 46.41 4.18 6 (69.79, 70.47) 0.68
leaky Rope-Longer 0.0001 68.74 4.59 56.80 5.87 4 (64.24, 73.24) 9.00
silu Rope-Longer 0.0001 67.31 3.54 51.40 8.62 5 (64.21, 70.41) 6.20
leaky Abs 0.0001 65.46 6.64 50.75 11.48 9 (61.12, 69.8) 8.68
silu Abs 0.0001 64.15 2.94 42.68 8.87 6 (61.8, 66.5) 4.70
leaky Abs 0.1000 69.20 4.78 54.43 8.87 6 (65.38, 73.02) 7.64
silu Abs 0.1000 65.30 8.34 52.33 10.38 9 (59.85, 70.75) 10.90
leaky RMHA-4 0.0001 68.65 0.48 43.24 0.35 6 (68.27, 69.03) 0.76
silu RMHA-4 0.0001 69.70 0.64 43.75 1.44 6 (69.19, 70.21) 1.02
leaky RMHA-4 0.1000 58.72 3.28 41.22 6.16 6 (56.1, 61.34) 5.24
silu RMHA-4 0.1000 60.77 1.59 40.89 5.90 6 (59.5, 62.04) 2.54
leaky RopeOne 0.0001 69.77 7.86 54.24 17.65 3 (60.88, 78.66) 17.78
silu RopeOne 0.0001 62.14 0.74 37.49 3.31 3 (61.3, 62.98) 1.68
leaky RopeOne 0.0001 67.82 6.59 54.10 12.14 6 (62.55, 73.09) 10.54
silu RopeOne 0.0001 66.24 4.18 49.30 11.44 6 (62.9, 69.58) 6.68
leaky Rope 0.0001 65.29 6.06 49.86 11.15 7 (60.8, 69.78) 8.98
silu Rope 0.0001 63.93 3.25 46.10 7.61 5 (61.08, 66.78) 5.70
leaky Rope 0.1000 62.76 9.47 49.77 12.13 6 (55.18, 70.34) 15.16
silu Rope 0.1000 61.24 8.88 46.52 12.55 6 (54.13, 68.35) 14.22
leaky Rotatory 0.0001 67.71 5.57 52.99 11.95 9 (64.07, 71.35) 7.28
silu Rotatory 0.0001 67.07 3.63 54.88 4.54 6 (64.17, 69.97) 5.80
leaky Rotatory 0.1000 68.07 4.68 56.31 6.21 6 (64.33, 71.81) 7.48
silu Rotatory 0.1000 63.07 8.92 49.38 12.30 10 (57.54, 68.6) 11.06
leaky None 0.0001 55.0 34.9 5
Table 9: Men Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
leaky None NaN 78.95 1.25 53.37 1.21 6 (77.95, 79.95) 2.00
silu None NaN 78.49 0.98 52.83 0.74 10 (77.88, 79.1) 1.22
leaky Learnt + Con NaN 78.34 1.28 54.14 1.39 6 (77.32, 79.36) 2.04
silu Learnt + Con NaN 79.32 1.50 53.68 1.97 6 (78.12, 80.52) 2.40
leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79 6 (80.18, 81.06) 0.88
silu Rotatory + Con. NaN 79.61 2.17 54.80 2.71 6 (77.87, 81.35) 3.48
leaky Abs + Con NaN 79.71 1.28 54.93 1.50 6 (78.69, 80.73) 2.04
silu Abs + Con NaN 79.34 1.67 54.23 1.84 10 (78.3, 80.38) 2.08
leaky Learnt NaN 78.64 2.05 54.30 2.41 6 (77.0, 80.28) 3.28
silu Learnt NaN 79.82 1.88 54.30 2.95 8 (78.52, 81.12) 2.60
leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93 6 (79.0, 81.42) 2.42
silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55 6 (79.91, 80.67) 0.76
leaky Rotatory-Longer NaN 78.00 1.10 52.30 1.05 6 (77.12, 78.88) 1.76
silu Rotatory-Longer NaN 76.41 1.00 51.48 1.10 6 (75.61, 77.21) 1.60
leaky Abs NaN 75.27 2.30 49.61 2.90 6 (73.43, 77.11) 3.68
silu Abs NaN 73.27 3.25 47.22 3.75 6 (70.67, 75.87) 5.20
leaky RMHA-4 NaN 78.85 0.51 53.32 0.44 12 (78.56, 79.14) 0.58
silu RMHA-4 NaN 76.59 0.78 51.54 0.43 6 (75.97, 77.21) 1.24
leaky RopeOne NaN 78.82 0.62 53.45 0.42 6 (78.32, 79.32) 1.00
silu RopeOne NaN 78.27 1.19 52.85 0.88 6 (77.32, 79.22) 1.90
leaky Rope NaN 78.97 0.33 53.64 0.27 6 (78.71, 79.23) 0.52
silu Rope NaN 77.56 1.17 52.04 1.25 6 (76.62, 78.5) 1.88
leaky Rotatory NaN 76.56 2.17 51.09 2.29 6 (74.82, 78.3) 3.48
silu Rotatory NaN 76.21 1.77 51.06 2.04 6 (74.79, 77.63) 2.84
leaky None 0.0001 78.2 57.3 5
Table 10: Games Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
A.2 Resources
Due to the instability of these models, we had to run many seeds to get some confident results. We started by running each one of the model configurations three times. After that, we decided to run more seeds based on the confidence of each one of the cases. We tried to keep at least a confidence range of ten points. We used $V100$ GPUs of 32GBs. For $men$ and $fashion$, based on the encoding, the training times go from $4$ to $6$ hours. For $Beauty$, they went from $22$ to $26$ hours. Based on the team’s resource ability, these experiments were carried out over $6$ months, lasting between $1$ and $8$.
We used the following seeds: $42$, $43$, $44$, $45$, $46$, $1809$, $1810$, $1811$, $1812$, $1741$, $1742$, $1743$, $1744$, $1745$, $123$, $234$, $345$, $456$, $567$, $678$, $789$, $890,102$, $203$, $304$, $405$, $506$, $607$, $708$, $808$, $901$, $1001$, $987$, $8765$, $7654$, and $6543$.
A.3 Hyperparameters
In this section, we present the hyperparameters used for training. As mentioned above, we use the same hyperparameters as CARCA Rashed et al. (2022). In some cases, we experimented with the upper bound from the norm in the encoding and the training length.
Hyperparameter Men Fashion Games Beauty
Learning Rate 0.000006 0.00001 0.0001 0.0001
Max Seq. Length 35 35 50 75
Number of attention blocks 3 3 3 3
Number of attention heads 3 3 3 1
Dropout Rate 0.3 0.3 0.5 0.5
L2 reg. weight 0.0001 0.0001 0 0.0001
Embedding Dimension d 390 390 90 90
Embedding Dimension g 1950 1950 450 450
Residual connection in the cross-attention block No No Yes Yes
Table 11: Hyperparameters with corresponding options were used in the original CARCA model. This parameter were written in tensor flow. In "L2 reg. weight", the value 00 corresponds to no bound, which in Pytorch is $Nan$, and we denote it as None .
A.4 Deviations over all the trainings
Here, we present the table with all the deviation values from the training. In the table in the main text, we considered those training where we use the default values for the upper bound, i.e., $0.0001$, and $None$ for the Games dataset. We can see that the trends pointed out in Section 5.1 are still valid in this table. The encodings with concatenation are more stable, and RMHA-4 is the most stable one. We can see that with longer trainings Rotatory encodings become more stable than Rope
encoding HIT dev NDCG dev runs CI-length
Abs 7.20 9.42 10.57 8.42
Abs + Con 4.67 6.58 8.00 6.36
Learnt 7.43 9.98 9.07 10.17
Learnt + Con 4.09 6.53 6.93 5.66
None 3.05 5.08 6.71 5.62
RMHA-4 1.37 2.64 6.29 1.98
Rope 4.03 6.55 5.14 6.52
Rotatory 3.92 6.10 9.00 4.92
Rotatory + Con. 3.16 5.12 6.71 4.47
RopeOne 2.82 6.54 4.29 5.46
Rope-Longer 3.33 8.15 5.25 5.99
Rotatory-Longer 1.31 1.35 4.00 3.00
Rotatory + Con.-Longer 0.99 1.24 6.00 1.59
RMHA-4-Longer 0.71 4.00 6.00 1.14
Table 12: Deviations among the different encodings. Those per head, like RMHA-4 or Rope , are more stable than others.
A.5 Losses
This section shows some loss cases from our trainings showing the stability of RMHA-4 . The first image shows four test losses for $0.0001$ as upper bound, with silu activation. The first row corresponds to the None encoding, while the second row is the RMHA-4 encoding. Each row contains the losses of four different seeds. The seeds are not paired vertically but randomly selected.
The second image shows the validation Losses for the analogous case in the Men Dataset.
Figure 2: Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ ad silu. Dataset: Fashion
Figure 3: Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ ad silu. Dataset: Men

### v2

@@
 Positional encoding is not the same as context: A study on positional encoding for sequential recommendation
-Abstract
-The expansion of streaming media and e-commerce has led to a boom in recommendation systems, including Sequential recommendation systems, which consider the user’s previous interactions with items.
-In recent years, research has focused on architectural improvements such as transformer blocks and feature extraction that can augment model information.
-Among these features are context and attributes.
-Of particular importance is the temporal footprint, which is often considered part of the context and seen in previous publications as interchangeable with positional information.
-Other publications use positional encodings with little attention to them.
-In this paper, we analyse positional encodings, showing that they provide relative information between items that are not inferable from the temporal footprint.
-Furthermore, we evaluate different encodings and how they affect metrics and stability using Amazon datasets.
-We added some new encodings to help with these problems along the way.
-We found that we can reach new state-of-the-art results by finding the correct positional encoding, but more importantly, certain encodings stabilise the training.
-1 Introduction
-Sequential recommendation systems (SRS) have been booming within referral systems in the last few years since most of these systems are based on previous user behaviour and are present in almost all e-commerce, social media, and streaming platforms.
-Their effectiveness has led to various definitions of sequential recommendations and to their division into branches.
-In general, sequential recommendation systems Fang et al.
-(2020), refer to recommendation systems (RS) where the interaction between the user and items is recorded sequentially.
-SRS are also related to session-based or session-aware recommendations.
-The latter two terms can be viewed as the sub-types of sequential recommendation when the interactions log is restricted to those made in the current session.
-In this work, we adhere to the notation used in Fang et al.
-(2020) and use the term object to summarise the interactions.
-"Behavior object" means what the user interacts with, identified by an ID or a group of items.
-It may include text, images, and time.
-We will use "item(s)" for convenience.
-The interaction between items and users can use many different types of features.
-A typical split of those is attributes and context, where the attributes are features provided by the item, like a text description or an image, and the context refers to those features related to the interaction itself, like the time that a user tends to buy or search some streaming content.
-A division based on the target task exists between the next-item recommendation and the next-basket recommendation Fang et al.
-(2020).
-In the next-item recommendation, a behaviour contains only one object (i.e., item), which could be a product, a song, a movie, or a location.
-In contrast, a behavior contains more than one object in next-basket recommendation.
-When dealing with sequential systems, it is essential to consider the sequential nature of the data.
-Therefore, initial deep learning techniques utilised Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN) to address this issue.
-However, attentional models have become increasingly popular due to their effectiveness, leading to multiple approaches utilising this technique Wang et al.
-(2021).
-Currently, most state-of-the-art models use attentional networks.
+Abstract.
+The rapid growth of streaming media and e-commerce has driven advancements in recommendation systems, particularly Sequential Recommendation Systems (SRS).
+These systems employ users’ interaction histories to predict future preferences.
+While recent research has focused on architectural innovations like transformer blocks and feature extraction, positional encodings, crucial for capturing temporal patterns, have received less attention.
+These encodings are often conflated with contextual, such as the temporal footprint, which previous works tend to treat as interchangeable with positional information.
+This paper highlights the critical distinction between temporal footprint and positional encodings, demonstrating that the latter offers unique relational cues between items, which the temporal footprint alone cannot provide.
+Through extensive experimentation on eight Amazon datasets and subsets, we assess the impact of various encodings on performance metrics and training stability.
+We introduce new positional encodings and investigate integration strategies that improve both metrics and stability, surpassing state-of-the-art results at the time of this work’s initial preprint.
+Importantly, we demonstrate that selecting the appropriate encoding is not only key to better performance but also essential for building robust, reliable SRS models.
+copyright: noneconference: ; ; price: doi: isbn:
+1. Introduction
+Sequential Recommendation Systems (SRS) have become integral to personalizing user experiences by predicting the next item of interest based on historical interaction sequences.
+These systems are widely deployed across e-commerce, social media, and streaming platforms and significantly enhance user engagement and satisfaction.
+As defined in (Fang et al., 2020), SRS records user-item interactions sequentially, closely related to session-based and session-aware recommendations, which focus specifically on interactions within a single session.
+Figure 1.
+Illustration of our proposed Rotatory  embeddings.
+Similar to RoPE , it encodes positional information through the angles of the embeddings.
+However, unlike RoPE , which assigns angles proportionally based on position, Rotatory learns these angles during training.
+It allows the model to determine the most relevant positional information and adjust their significance accordingly.
+Capturing the order of data is essential in sequential systems.
+Initially, Deep Learning techniques utilized Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN) to model sequential dependencies.
+However, attentional-based models have gained prominence due to their superior effectiveness, leading to numerous self-attention approaches (Wang et al., 2021).
+Most state-of-the-art (SOTA) models in SRS now leverage attentional-based networks.
+A critical component of these systems is positional encoding (PE), which imparts information about the order of items within a sequence.
+The choice and implementation of PEs are crucial for optimizing performance and stability.
+Despite the widespread adoption of various PE techniques, there is a noticeable gap in research concerning the stability of model performance across different runs.
+However, many existing publications focus on enhancing performance metrics without adequately addressing the underlying stability issues introduced by different encoding strategies.
+Other studies underestimate the relevance of positional encoding, considering them part of the context.
+To bridge this gap, we aim to investigate the stability problem in SRS models through positional encodings.
+We conduct extensive experiments to evaluate how different positional encodings influence these models’ performance and stability.
+Building on this foundation, we introduce a novel positional encoding method termed Rotatory , along with its concatenated variant, Rotatory + Con , and ablations of the previous encoding method.
 To summarise, our contribution is fourfold:
-  1. 1.
-    We analyse existing work and highlight that more than the standard usage of $5$ runs is needed to ensure the reliability of the results.
-  2. 2.
-    We propose new encodings. Based on our first experiments with existing encodings, we propose new encodings to improve metrics and stability.
-  3. 3.
-    We perform a detailed analysis of the types of encodings, their features (in-head, affected heads, etc.) and their ablations to see which factors improve the stability and the metrics.
-  4. 4.
-    We found correlations between the sparsity of the dataset and the encodings to get the best results.
-This precise data dependence, together with our study, provides a road map for the usage of encodings for SRS with transformer architectures.
-2 Related Work
-From an architectural perspective, SRS has evolved over the last two decades.
-Traditional methods in the field applied classic ML techniques, such as KNN for item-based systems (Davidson et al.
-(2010), Linden et al.
-(2003)), or applications of Markov Chains for different orders (He et al.
-(2017), He et al.
-(2016b), He and McAuley (2016a)).
-A common approach due to the sparsity of the interactions in RS is those based on matrix factorisation like DeepFM Guo et al.
-(2017) or CFM Xin et al.
-(2019).
-Among the different DL architectures, RNNs can capture sequence information and represent items based on their position in the list.
-One of the earliest and best known was Cho et al.
-(2014).
-The first approach from DL came using RNN, the most common architecture then.
-We can find models such as GRU4Rec Hidasi et al.
-(2016) or DREAM Yu et al.
-(2016).
-Another of the first approaches via DL was based on CNN models that can capture local features and consider time information in the input layer (Tang and Wang (2018), Tuan and Phuong (2017)).
-Among these were Caesar Li et al.
-(2021) and NextItNet Yuan et al.
-(2018).
-One of the first approaches using attention models was NARM Li et al.
-(2017), an encoder-decoder framework that uses attention at the bottleneck.
-The introduction of transformers and self-attention in sequential models came from AttRec Zhang et al.
-(2018).
-Over time, different types of information have been identified to predict what a user might buy next.
-There are three main types of models for this: attribute-aware Rashed et al.
-(2022), context-aware Zhou et al.
-(2020), and time-aware Li et al.
-(2020).
-Attribute-aware models focus on the information provided by the items themselves.
-Context-aware models take into account the characteristics of the user’s interaction.
-Time-aware models are a subgroup of context-aware models that use the timing of the user’s interactions.
-Using transformers in architecture has led to the introduction of positional embeddings.
-By default, the attention heads work like bags of tokens, and the positional information needs to be added to understand the sequential order of the tokens.
-Some approaches use time-aware information to replace the position, like in the case of CARCA Rashed et al.
-(2022).
-While this approach achieved excellent results, it faces some limitations.
-Replacing the positional embeddings with time stamps assumes that the order of the items will be implicitly deduced from this stamp.
-This can be more challenging to the model than simply an explicit representation.
-Furthermore, recent research in NLP Chen et al.
-(2021) has shown that introducing positional information within the attention, instead of the first head input, produces better results, which can not be done through the time stamps.
-One of the first models considering the sequential information was GRU4Rec Hidasi et al.
-(2016), which concentrates on extracting information from the sequence of items instead of the user being an anonymous user model.
-A successful model from the context-aware point of view was DeepFM Guo et al.
-(2017), which extracted through the combination of factorisation models and deep learning extracts information from the context.
-Like GRU4Rec, using a self-attention layer to balance short-term intent and long-term preference can be found in SASRec Kang and McAuley (2018a).
-The extraction through a sequence of transformer encoders produces an output, which is compared with the target items through dot-product.
-This dot-product approach combines information from past behaviour and target representation, which was improved by an encoder-decoder architecture, where the encoder extracts past information, and the decoder combines the embedding representation for the target.
-This produced one of the SOTA models, CARCA Rashed et al.
-(2022).
-Another path of improvement from SASRec came from training the bidirectional transformer to infer the item-item relationship from the user’s historical interactions using the Cloze task in BERT4Rec Sun et al.
-(2019).
-Another one was TiSASRec Li et al.
-(2020), which further improves SASRec by considering time intervals between items.
-We will focus on the position encoding applied to the most common datasets among these papers.
-For a deeper review of the different encoding techniques, one can see Zhao et al.
-(2024).
-Since there is no systematic study, each work addresses the encoding differently.
-In AttRec Zhang et al.
-(2018), they "furnish the query and key with time information by positional embeddings" using absolute embeddings.
-SASRec used absolute learnable position embeddings in the input sequence.
-They also tried the fixed position embedding as used in Vaswani et al.
-(2017) but found that this led to worse performance.
-In CARCA, which will be our base model, they tried using absolute encodings again, giving even worse results than no positional encoding.
-This disparity of results merits a more detailed analysis as in the present work.
-3 Positional encodings
+  1. (1)
+    We analyze existing work and highlight that more than the standard usage of $5$ runs is needed to ensure the reliability of results.
+  2. (2)
+    We conduct a detailed analysis of various encoding types, their features (such as in-head and affected heads), and perform ablation studies to identify factors that improve stability and performance.
+  3. (3)
+    We introduce a novel rotatory-based encoding method, alongside integration techniques, such as concatenation, which have not been explored previously.
+These contributions aim to improve both performance metrics and training stability.
+  4. (4)
+    Our findings show that encoding effectiveness is primarily influenced by dataset deviation, rather than sparsity or data type.
+This insight, combined with our analysis, offers a clearer guide for using encodings in SRS with transformer architectures.
+2. Related Work
 Encoding Learnable Type Range Layers Concatenation Integration
@@
 Abs + Con Yes APE Fix. Vector Encoding Yes Joint
-Rotatory No APE Unl. Vector Encoding No Multi.
-Rotatory + Con. Yes APE Unl. Vector Encoding Yes Joint
+Rotatory Yes APE Unl. Vector Encoding No Add
+Rotatory + Con Yes APE Unl. Vector Encoding Yes Joint
 RMHA-4 No RPE 4 All No Multi.
@@
 None No N/A N/A 0 N/A N/A
-Table 1: Encoding types: ’Type’ refers to absolute and relative.
-’Learnable’ indicates whether the encoding parameters are learned during training.
-We also consider learnable to have a concatenation since there are learnable parameters.
-’Range’ specifies the positions the encoding covers: ’Fix.’ fixed by the given length, ’Unl.’ means unlimited, while $4$ refers to the relative range.
-’Layers’ represents the number of layers the encoding uses: "Vector Encoding" means they are added at the start by vectors (left in Fig: 1), All means in all the transformer blocks and First, at the first encoding block.
-’Concatenation’ indicates whether there is a concatenation.
-’Integration’ describes how the encoding is integrated with the model architecture: ’Add’ indicates a vector sum, ’Joint’ the concatenation, and ’Multi.’ for multiplication of matrices for those in the heads or element-wise for the others.
-Finally, None refers to no encoding.
-Figure 1: APE vs RPE : While APE encodings introduce the positional information as vectors before the $V$, $Q$ and $K$, projections, RPE add this information at the coefficient level in $K$ and $Q$.
+Table 1.
+Encoding types: ’Type’ specifies absolute or relative.
+’Learnable’ indicates if encoding parameters are trained, including concatenation for learnable encodings.
+’Range’ denotes position coverage: ’Fix.’ for fixed length, ’Unl.’ for unlimited, and $4$ for relative range.
+’Layers’ refers to encoding application: ’Vector Encoding’ adds vectors at the start (left in Fig 3), ’All’ applies to all transformer blocks, and ’First’ to the first block.
+’Concatenation’ shows if concatenation is used.
+’Integration’ describes the method: ’Add’ for vector sum, ’Joint’ for concatenation, and ’Multi.’ for matrix multiplication in heads or element-wise otherwise.
+’None ’ means no encoding.
+2.1. Architectures & Encoding in SRS
+From an architectural perspective, SRS have significantly evolved over the past two decades.
+Early approaches employed K-Nearest Neighbors (KNN) for item-based recommendations (Davidson et al., 2010; Linden et al., 2003) and Markov Chains for modeling sequential interactions (He et al., 2017, 2016b; He and McAuley, 2016a).
+To address interaction sparsity, matrix factorization methods like DeepFM (Guo et al., 2017) and CFM (Xin et al., 2019) were introduced, enhancing the ability to handle sparse data effectively.
+With the advent of Deep Learning, RNNs became popular for capturing sequence information and item positions, exemplified by models such as GRU4Rec (Hidasi et al., 2016) and DREAM (Yu et al., 2016).
+Concurrently, CNNs were utilized to capture local features and temporal information, as seen in Caesar (Li et al., 2021) and NextItNet (Yuan et al., 2018), among others (Tang and Wang, 2018; Tuan and Phuong, 2017).
+Attention mechanisms were first introduced in NARM (Li et al., 2017), and transformers with self-attention were pioneered in SRS by AttRec (Zhang et al., 2018).
+Different types of models are used to predict user behavior: attribute-aware (Rashed et al., 2022), context-aware (Zhou et al., 2020), and time-aware (Li et al., 2020).
+These models focus on item information, user interaction characteristics, and the timing of interactions, respectively.
+The integration of transformers has introduced positional embeddings.
+Some approaches, like CARCA (Rashed et al., 2022), use time-aware information to replace traditional positional encodings, assuming that the order of items can be deduced from timestamps.
+However, recent NLP research (Chen et al., 2021) has shown that incorporating positional information directly within the attention mechanism, rather than relying on timestamps, yields better results.
+One of the initial models to capture sequential information was GRU4Rec (Hidasi et al., 2016), which extracts data from item sequences.
+The context-aware DeepFM (Guo et al., 2017) combines factorization methods with deep learning to leverage contextual information.
+Similarly, SASRec (Kang and McAuley, 2018a) employs a self-attention layer to balance short-term intent and long-term preferences, using transformer encoders to compare past behavior with target items via dot-products.
+This encoder-decoder architecture, which extracts past information while synthesizing target embeddings, led to the development of CARCA (Rashed et al., 2022).
+We focus on positional encoding across various datasets used in these studies.
+Each work approaches encoding differently: AttRec (Zhang et al., 2018) employs absolute embeddings to incorporate time information into the query and key, while SASRec utilizes absolute learnable position embeddings in the input sequence.
+CARCA experimented with absolute encodings but observed worse performance compared to using no positional encoding.
+Recent advancements, such as EulerFormer (Tian et al., 2024), introduce complex vector attention to integrate semantic and positional information.
+This variety in encoding strategies highlights the need for a systematic analysis, which we provide in this work.
+Figure 2.
+(1) Vector encoding added before Transformer blocks, (2) First head RPE: Relative encoding applied only to the first block (RopeOne ), and (3) All head RPE: Relative encoding integrated into every block (RMHA-4 ).
+Figure 3.
+APE vs RPE : While APE encodings introduce the positional information as vectors before the $V$, $Q$ and $K$.
+RPE add this information at the coefficient level to $K$ and $Q$.
 Parts affected by the position information appear in red.
-3.1 Positional encodings
-This section introduces the different types of encoding techniques used in this work.
-For a deeper review of the various encoding methods, check Zhao et al.
-(2024).
-Following Zhao et al.
-(2024), there are two prominent families of encoding methods.
-First, we have absolute positional encoding (APE ).
-These encodings are added before the first Transformer block, and the positional information provided is absolute, from the initial token to the last one.
-Second, there are the relative encoding methods (RPE ), which provide information about those items around the target item.
-These encodings are implemented inside the transformer’s head, more precisely in the Queries and Keys.
-In contrast to the common usage of encoding in Transformers for NLP tasks in RS, the encoding is reversed; the first item in the encoding order is the last one, and padding is applied on the left side Kang and McAuley (2018a).
-We chose this approach because the most relevant items are usually the ones that were recently visited or clicked.
-Due to the sinusoidal functions used in APE , the most pertinent items might be at the start because the assigned frequency becomes less distinguishable the later an item appears.
-We did some preliminary experiments to check if this is also right for another encoding, and the reversed encoding always produced better results, so in what follows, all the encoding will be considered as reversed.
+2.2. Positional encodings
+Positional encoding enables transformer-based models to incorporate item order within sequences.
+Encoding methods can be categorized into absolute and relative encodings (Zhao et al., 2024) (Figure 3).
+In Table 1, we have a detailed classification of the main encodings in this paper.
+Absolute Positional Encoding (APE )
+They assign a unique position vector to each item based on its position in the sequence.
+These encodings are added before the first Transformer block, providing absolute positional information.
+Models like SASRec (Kang and McAuley, 2018a) utilize learnable absolute position embeddings, allowing the model to learn position representations during training.
+There are two common variations of absolute positional encodings: fixed (Abs ) and learnable (Learnt ).
+Absolute Positional Encoding (Abs )
+Absolute positional encoding adds fixed values to item embeddings based on their absolute position in the sequence.
+The most common approach uses sinusoidal functions, as seen in BERT (Devlin et al., 2019).
+This encoding is added at the input layer as part of the item encoding (Figure, left 3).
+The encoding for an item at position $pos$ in a sequence is calculated using sine and cosine functions:
+(1) $PE_{(pos,2i)}=\sin\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$
+(2) $PE_{(pos,2i+1)}=\cos\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$
+where $i$ is the dimension, and $d$ is the embedding dimension. This encoding is added to the original embedding:
+(3) $x=x+PE_{(pos,*)}$
+Learnable Encoding (Learnt )
+Learnable position embeddings treat positional information as trainable parameters.
+Unlike fixed encodings, the model learns optimal position representations based on the data context, denoted as learnable encoding (Learnt ).
+Positional encodings are added to the input sequence embeddings.
+Relative Positional Encoding (RPE )
+Relative Positional Encoding (Figure 3, right) encodes the relative distances between items instead of their absolute positions, enabling models to better capture contextual relationships within user interactions.
+Unlike Absolute Positional Encoding (APE ), which requires fixed input lengths and provides positional information from the start to the end of the sequence, RPE offers flexibility by focusing on the relationships between items regardless of their absolute positions.
+A downside of these positional encoding mechanisms is that they are significantly slower during training and inference since they add an extra step in the self-attention layer (Press et al., 2022).
 Self-attention
-enables models to weigh different parts of input sequences differently, allowing them to focus on relevant information during processing.
-The self-attention mechanism was introduced in Vaswani et al.
-(2017), which laid the foundation for the transformer architecture.
-In self-attention, the input sequence is typically represented as a set of vectors, where each vector corresponds to a position in the sequence.
-This allows each position to attend to all other positions and captures dependencies and relationships within the sequence.
-For a given dimensionality $d$ of the item representation and a hidden dimension $H<d$, the attention weight assigned to each position is computed based on the similarity (dot product) between the query $Q\in\mathbb{R}^{d\times H}$, key $K\in\mathbb{R}^{d\times H}$, and value $V\in\mathbb{R}^{d\times H}$ matrices associated with the positions.
-These learnable matrices are derived from linear transformations of the original input $X\in\mathbb{R}^{n\times d}$ by multiplication, where $n$ is the number of items in the sequence.
-The attention weight ($\alpha_{ij}$) assigned to position $i$ for attending to position $j$ is computed using the following formula:
-$\alpha_{ij}=\frac{{\exp((X_{i}\,Q)\cdot(X_{j}\,K)^{T}/\sqrt{d})}}{{\sum_{t=1}^{n}\exp((X_{i}\,Q)\cdot(X_{t}\,K)^{T}/\sqrt{d})}}$ (1)
-Here, $Q_{i}$ and $K_{j}$ represent the query and key matrices associated with positions $i$ and $j$, respectively.
-The dot product is scaled by $\sqrt{d}$ to prevent vanishing gradient problems.
-The resulting attention weights are then used to compute a weighted sum of the value matrices ($V_{j}$) associated with all positions:
-$\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}X\cdot V_{j}$ (2)
-This weighted sum, representing the attended information, is then used as the output of the self-attention mechanism. The process is repeated for all positions in the sequence.
-Absolute Positional Encoding (APE )
-Absolute positional encoding adds fixed values to the embeddings of items based on their absolute position in the sequence.
-This means all the items have the same initial token as a reference.
-The most common approach is based on trigonometry, like in BERT Devlin et al.
-(2019), the so-called sinusoidal encoding.
-Absolute encoding is added at the input layer as part of the item encoding, 1.
-The most common practice is to add a vector encoding the position.
-The absolute positional encoding for a token at position $i$ in a sequence of length $N$ can be calculated using trigonometric functions such as sine and cosine. A common formula is:
-$PE_{(pos,2i)}=\sin\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$ (3)
-$PE_{(pos,2i+1)}=\cos\left(\frac{{pos}}{{10000^{(2i/d)}}}\right)$ (4)
-where $pos$ is the position of the item, $i$ is the dimension, and $d$ is the dimension of the embedding.
-This encoding is usually added to the original embedding, and we will refer to it along the text as absolute encoding (AE).
-For a given item embedding $x$ at position $pos$, we replace it as follows:
-$x=x+PE_{(pos,*)}$ (5)
-Another way of making the positional information more explicit and separated from the rest of the context and attributes is by concatenation, where we divide the dimension of the embedding into the context and attribute information and the positional encoding information:
-$x=Concat\,(x,PE_{(pos,*)})$ (6)
-We apply a Linear layer right after to recover the original dimension and do not change the number of parameters in the model.
-We will refer to them as concat absolute encoding (CAE).
-We added the position information orthogonally to make the model gradually rely on the position, producing more stable results, Section 5.
-Learnable Encoding
-Learnable position embeddings take a different approach by treating the positional information as parameters that can be updated during training.
-Instead of relying on predefined patterns, the model learns the optimal representation of positions based on the data context.
-We will denote these embeddings as learnable encoding (LE).
-In this case, we add a positional encoding, which is merely a learnable vector.
-We also apply a concatenation approach denoting as concatenated learnable encoding (CLE).
-In this case, their results will be generally worse than the classic residual connection.
-Relative Positional Encoding (RPE )
-A disadvantage of using APE is that it requires a fixed length of the input sequence and does not directly capture the relative positions of one item to another, whether the other is further ahead or backward.
-Relative position encoding (RPE ) was proposed to solve these issues, 1.
-Shaw et al.
-(2018) uses relative position encoding (RPE ) instead of APE and adds the position embeddings to the key and optional value projections instead of the input.
-A simplified version deleting the position embeddings in value projections in TranformersXL Dai et al.
-(2019) showed better performance on language modelling tasks.
-The length of the user’s session or action log varies significantly from user to user, suggesting that these encoding techniques may be of value to SRS.
-The relative positional encoding can be computed using trainable parameters.
-The specific formula may depend on the implementation; in our case, we follow the one in Shaw et al.
-(2018), where the position is given inside the attention.
-The relation between two inputs $X_{i}$ and $X_{j}$ is given by $a_{ij}^{V},a_{ij}^{K}\in\mathbb{R}^{d_{a}}$, for values and keys.
-This information can be shared among heads, but we consider the case in which information is not shared.
-For the relative case, the attention weight ($\alpha_{ij}$) can be written as:
-$\frac{{\exp((X_{i}\,Q)\cdot(X_{j}\,K+a_{ij}^{K})^{T}/\sqrt{d})}}{{\sum_{t=1}^{n}\exp((X_{i}\,Q)\cdot(X_{j}\,K+a_{ij}^{K})^{T}/\sqrt{d})}}$ (7)
-We also introduce the position at the value product as follows:
-$\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}(X\cdot V_{j}+a_{ij}^{V})$ (8)
-Relative positional encoding can be especially useful when dealing with long sequences, as it allows the model to generalise better to different distances between items and in cases where the local position information is more critical.
-As a drawback, since it has to be applied to each attention head, it tends to be slower Shaw et al.
-(2018).
-We will denote these embeddings as RMHA-4 , where we bound the relative indexes to a distance of $4$.
-We do not apply a concatenated version in this case.
-Rotary Encoding (Rope and Rotatory )
-Unlike absolute or relative positional encoding, which focuses on representing positions in a fixed coordinate system or relative to other items in a sequence, rotary encoding leverages rotational transformations to encode position information.
-"Rotary Positional Encoding" typically refers to a specific technique used in positional encoding, Su et al.
-(2024).
-We will use the notation Rope to refer to them.
-The motivation behind Rope is that the orientation or rotation of positional embeddings can provide valuable information about the relationship between items in a sequence.
-This approach is particularly relevant in scenarios where the inherent order or arrangement of items is not straightforward and can benefit from a more flexible representation, like a user searching for more than one item in the same session.
-These embeddings capture both linear order and rotational relationships within the sequence.
-Rotary position encodings apply a rotation to the Query and Key matrices, which preserves the dot product among the vectors.
-In the two dimensional case for a given input vector $x$ the encoding function $f$ for given angle $\theta$ is defined as follows:
-$f_{\theta}(\mathbf{x})=\mathbf{R}(\theta)\cdot\mathbf{W}\cdot\mathbf{x}$ (9)
-Here, $\mathbf{R}(\theta)$ represents the rotation matrix parameterized by an angle $\theta$, expressed using sines and cosines, and $\mathbf{W}$ is either $K$ or $Q$.
-More explicit, for the $K$ case we have:
-$\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\cdot\begin{bmatrix}K_{11}&K_{12}\\ K_{21}&K_{22}\end{bmatrix}\cdot\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$ (10)
-The straightforward generalisation of this to higher dimensionality divides the vector dimension into two-dimensional chunks and applies the rotation to each of them.
-Since this implementation is costly, the actual implementation expresses this rotation as two vector multiplications and one vector addition, as RoFormer Su et al.
-(2024).
-Since we want to analyse the effect of the rotational position for sinusoidal, we propose here a new encoding, Rotatory , which is concatenated version Rotatory + Con.
-.
-Instead of directly applying sinusoidal functions, rotary encodings introduce an additional rotation mechanism by separating the angles into sine and cosine components and then alternating them across the embedding dimension.
-By alternating sine and cosine embeddings, the model can capture different phases of positional information, effectively introducing a rotational component to the encoding.
-This rotation helps mitigate potential issues, such as vanishing gradients that can arise with traditional positional encoding, which is especially important for stability.
-So, while the term "rotary" might not directly imply physical rotation, it refers to the rotational mechanism introduced in the encoding process to enhance the representation of positional information in Transformer models.
-The implementation is similar to that from Rope , but applying the rotation matrices to the vector encodings added to the input items.
-Dataset Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg
+This mechanism allows models to weigh different parts of the input sequence differently, focusing on relevant information during processing.
+Introduced in (Vaswani et al., 2017), the self-attention mechanism computes attention weights based on the similarity between queries ($Q$), keys ($K$), and values ($V$):
+(4) $\alpha_{ij}=\frac{\exp\left((Q_{i}K_{j}^{T})/\sqrt{d}\right)}{\sum_{t=1}^{n}\exp\left((Q_{i}K_{t}^{T})/\sqrt{d}\right)}$
+(5) $\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}V_{j}$
+Relative Positional Encoding Implementation (RMHA-4 )
+This encoding addresses the limitations of APE by encoding the relative positions within the attention mechanism.
+Models like NARM (Li et al., 2017) and TransformerXL (Dai et al., 2019) adopt RPE to enhance the modeling of contextual dependencies.
+In RPE , positional information is incorporated into the key and value projections, allowing the model to focus on the relationships between items dynamically.
+The attention weight with RPE is computed as:
+(6) $\alpha_{ij}=\frac{\exp\left((Q_{i}(K_{j}+a_{ij}^{K})^{T})/\sqrt{d}\right)}{\sum_{t=1}^{n}\exp\left((Q_{i}(K_{t}+a_{it}^{K})^{T})/\sqrt{d}\right)}$
+Additionally, positional information can be added to the value projections:
+(7) $\text{Attention}(Q,K,V)=\sum_{j=1}^{N}\alpha_{ij}(V_{j}+a_{ij}^{V})$
+We will consider the case which bounds the relative indexes to a distance of $4$ and called it RMHA-4 , Relative Multi-Head Attention with distance $4$.
+RoPE (RoPE )
+Rotary Positional Encoding (Su et al., 2024) encodes positional information by applying rotational transformations to the Query and Key matrices.
+This preserves the dot product between vectors, maintaining effective attention mechanisms.
+RoPE is particularly useful for complex item orders, providing a more flexible representation than traditional encodings.
+For a two-dimensional vector $\mathbf{x}$ and angle $\theta$:
+(8) $f_{\theta}(\mathbf{x})=\mathbf{R}(\theta)\mathbf{W}\mathbf{x}$
+where $\mathbf{R}(\theta)$ is a rotation matrix defined by sines and cosines, and $\mathbf{W}$ represents either $K$ or $Q$.
+Higher-dimensional vectors are split into two-dimensional chunks, with rotations applied to each.
+To reduce computational costs, RoFormer (Su et al., 2024) implements these rotations using two vector multiplications and one vector addition.
+3. Methodology
+In this section, we present our contributions to positional encodings by introducing variants of existing methods and new encodings.
+3.1. Encoding Variants
+The most commonly used encodings are Abs encodings, which are vector encodings added to the initial item vector by addition.
+This approach preserves the original embedding dimension without requiring additional encoding dimensions (Table 1).
+In contrast, non-vector encodings, such as Relative Positional Encoding (RPE) and Rotary Positional Encoding (RoPE), are typically introduced as multiplications within the attention mechanism.
+These general rules apply depending on whether encodings are relative or absolute.
+To further explore encoding strategies, we introduce an encoding variant that employs concatenation with a linear layer and activation function to revert to the original embedding dimension.
+To the best of our knowledge, this approach has not been previously utilized in SRS and has seen limited application in NLP.
+While concatenation benefits include separating semantic and positional information, it introduces a size increase in the architecture, which is particularly critical for Large Language Models.
+To mitigate this, we incorporate an additional layer, allowing for a gradual introduction of positional information without excessively increasing model complexity.
+Concatenated Absolute Encoding
+To enhance positional information, we introduce concatenated absolute encoding, which concatenates the positional encoding to the item embedding:
+(9) $x=\text{Concat}(x,PE_{(pos,*)})$
+A Linear layer is then applied to restore the original embedding dimension without increasing model parameters.
+This method explicitly incorporates positional information, allowing the model to gradually rely on positional cues and achieving more stable performance (Section 5).
+Concatenated Learnable Encoding (Learnt + Con )
+Additionally, we apply a concatenation approach to learnable encodings, denoted as concatenated learnable encoding (Learnt + Con ).
+However, Learnt + Con generally yields worse results compared to the classic residual connection (Section 5), likely because concatenation adds complexity in this case, making it harder for the model to effectively learn embeddings.
+Furthermore, the gradual introduction of positional information is already inherent in the learnable encoding mechanism.
+We do not apply concatenated versions for RoPE and RMHA-4 , as they are relative encodings.
+RoPE One Block (RopeOne )
+In our experiments (Section 5) we found that RMHA-4 produce more stable results than other encodings.
+To investigate whether the stability improvements of RMHA-4 were due to introducing positional information at each Transformer block or its role as an RPE -based head, we tested RoPE .
+The results showed that that RoPE was in a middle point between relative and absolute encodings regarding stability.
+We hypothesize that this intermediate stability is due to the application of RoPE across all Transformer blocks (($3$) in Figure 2).
+To investigate this further, we introduced RopeOne , which applies RoPE exclusively in the first Transformer block (($2$) in Figure 2).
+As detailed in Section 5, this modification yields slightly improved results.
+This implies that the rotary aspect, rather than merely being relative, is key to the observed improvements and leads us to the new encoding introduced below.
+3.2. Rotatory Encoding
+We propose a new encoding method termed Rotatory (Figure 1), with its concatenated variant referred to as Rotatory + Con This approach applies a rotation to define the position of each item.
+Unlike traditional methods that add positional information during self-attention, our method incorporates these rotations at the initial item embeddings, similar to sinusoidal encoding (Abs ).
+Vector embeddings offer the advantage of faster training and inference (Press et al., 2022).
+Therefore, we adopt a formulation similar to RoPE , applying the rotations at the outset.
+We enhance flexibility by making the rotation angles learnable, allowing the model to capture both absolute and relative positional relationships effectively.
+The positional encoding for an item at position $pos$ and embedding dimension $i$ is calculated using a trigonometric approach, modified to include alternating sine signs for a rotational effect.
+The encoding is defined as follows:
+(10) $PE_{(pos,2i)}=(-1)^{i}\sin\left(\theta_{(pos,i)}\right)$
+(11) $PE_{(pos,2i+1)}=\cos\left(\theta_{(pos,i)}\right)$
+Here, $pos$ is the position of the item in the sequence, $i$ represents the embedding dimension, and $\theta_{(pos,i)}$ is a learnable angle defined as:
+(12) $\theta_{(pos,i)}=\frac{E_{pos}}{10000^{(2i/d)}}\cdot 2\pi,\quad\forall pos\in[0,L),\forall i\in[0,H)$
+where $E_{pos}\in\mathbb{R}^{L\times H}$ are learnable embeddings.
+The term $(-1)^{i}$ in the sine function alternates the sign, introducing the rotation effect in the encoding.
+These positional encodings are then added to the original item embeddings by addition:
+(13) $x^{\prime}_{(pos)}=x_{(pos)}+PE_{(pos)}$
+where $x_{(pos)}$ is the original semantic embedding of the item at position $pos$, and $PE_{(pos)}$ is the learnable Rotatory  encoding.
+In the case of Rotatory + Con , the positional encoding is concatenated with the original item embedding and passed through a learned linear transformation for further adaptation.
+This learnability introduces flexibility, enabling the model to capture both absolute and relative positional information in a dynamic manner, which leads to improved performance in handling complex sequence relationships.
+4. Experimental Setup
+4.1. Datasets
+We selected several Amazon datasets to evaluate performance of various encoding types.
+For this purpose, we utilised four distinct real-world datasets extracted from product reviews on Amazon.com.
+These datasets are widely used for SRS under the leave-one-out protocol (He and McAuley, 2016b; Kang and McAuley, 2018b; Hou et al., 2019; Wu et al., 2020; Steck, 2019; Zhou et al., 2020).
+Since multiple versions of these datasets exist, we opted for the preprocessed versions provided by (Rashed et al., 2022).
+Dataset Users Items Interactions Attributes
+Men 34,244 110,636 254,870 2,048
+Fashion 45,184 166,270 358,003 2,048
+Games 31,013 23,715 287,107 506
+Beauty 52,204 57,289 394,908 6,507
+Table 2. Dataset Statistics.
+Dataset Avg Dev Hit Avg Dev NDCG CI-length Density
+Men 3.9143 8.2742 5.6686 $6.727\cdot 10^{-5}$
+Fashion 4.87 9.7814 9.2257 $4.765\cdot 10^{-5}$
+Games 1.124 0.952 1.616 $3.904\cdot 10^{-4}$
+Beauty 0.935 0.87 1.22 $1.320\cdot 10^{-4}$
+Table 3. Deviations among the different datasets without Encoding. This pattern is generally preserved along the different encodings. The density refers to the inverse of Sparsity.
+Covering various product categories, these datasets offer a broad source for training and evaluating recommendation models.
+They include essential features for each interaction, such as user IDs, item IDs, timestamps, and additional vector-based contextual information.
+As shown in Tables 2 and 3, they exhibit diversity in terms of sparsity, defined as the ratio of missing user-item interactions.
+This will be a crucial aspect of our analysis, along with the discreteness of the features.
+1.Beauty
+(Wu et al., 2020; Zhou et al., 2020; Kang and McAuley, 2018b) The dataset includes discrete and categorical attributes for all beauty products, including fine-grained categories and brands.
+Mainly categorical and discrete features.
+2.Video Games
+(Wu et al., 2020; Kang and McAuley, 2018b) The Video Games sub-dataset includes user interactions, reviews, and product details specific to the video game category, such as price, brand and categorical features.
+Most of the attributes here are discrete and categorical.
+3.Men
+(He and McAuley, 2016b) The men’s dataset encompasses a comprehensive collection of items falling under men’s clothing.
+The attributes are dense vectors from image-based features extracted from the last layer of a ResNet50 (Deng et al., 2009) on the ImageNet dataset (He et al., 2016a).
+4.Fashion
+(He and McAuley, 2016b; Hou et al., 2019) It contains six categories for men’s and women’s clothing. The features were extracted as dense using the same ResNet50 approach.
+4.2. Models and Training
+To implement CARCA, we converted the original TensorFlow model into PyTorch [^1] https://github.com/researcher1741/Position_encoding_SRS.
+We used the loss implementation from (Papso, 2023).
+The original TensorFlow version built upon prior models within the same framework.
+During our evaluation, we observed that the implementations checked metrics every $20$ epochs, assuming that the optimal performance point would be reached later.
+This approach was likely adopted due to the extensive number of epochs, exceeding $1000$, which increases the computational cost of evaluations.
+To achieve more reliable evaluations, we implemented a logarithmic measurement of the metrics, which resulted in slightly improved outcomes in certain cases.
+As in previous works (Rashed et al., 2022; Kang and McAuley, 2018a), we use an ADAM optimiser to minimise the binary cross-entropy loss of the CARCA model while masking the padded items to prevent them from contributing to the loss function.
+For a given sequence of items interactions for a user $u$ as $\{i^{u}_{1},\ldots,i^{u}_{N}\}$, we create an input list as $I^{u+}=\{i^{u}_{1},\ldots,i^{u}_{N-1}\}$ by deleting the last item of the list, a positive target list as $T^{u+}=\{i^{u}_{2},\ldots,i^{u}_{N}\}$ by shifting the input list by one, and a negative target list as $T^{u-}=\{i^{rand_{1}},\ldots,i^{rand_{N-1}}\}$ generated by random negative items.
+Then the loss is given by:
+$-\sum_{u\in U}\sum_{r\in T^{u+}\cup T^{u-}}\left[Y_{r}\log(\hat{Y}_{r})+(1-Y_{r})\log(1-\hat{Y}_{r})\right]$
+Here, $U$ represents the set of users, $T^{u+}\cup T^{u-}$ denotes the union of positive and negative items for user $u$, $Y_{r}$ is the observed interaction label for item $r$, and $\hat{Y}_{r}$ is the predicted probability of interaction with item $r$.
+Encoding Avg Dev Hit Avg Dev NDCG runs CI-length
+Abs 5.43 8.25 9.38 6.71
+Abs + Con 3.52 5.56 7.12 5.35
+Learnt 4.79 7.64 7.88 7.05
+Learnt + Con 3.17 6.36 6.62 4.42
+None 2.35 5.26 8.00 3.29
+RMHA-4 0.51 1.08 6.00 0.88
+RoPE 2.38 5.45 5.00 4.04
+Rotatory 2.46 5.01 8.38 3.48
+Rotatory + Con 2.48 5.18 6.75 3.49
+RopeOne 2.82 6.54 4.29 5.46
+RoPE-Longer 3.33 8.15 5.25 5.99
+Rotatory-Longer 1.05 1.04 4.67 2.08
+Rotatory + Con-Longer 0.99 1.24 6.00 1.59
+RMHA-4-Longer 0.71 4.00 6.00 1.14
+Table 4.
+Deviations among the different encodings for the default upper bound values ($0.0001$ for all except NaN for Games).
+Relative encoding in all the heads, RMHA-4 , presents more stability, followed by our Rotatory versions.
+Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
 Beauty silu Rotatory-Longer 0.0001 61.87 1.04 42.60 0.98
 Beauty silu Rotatory 0.0001 61.72 1.14 42.54 1.52
-Beauty silu Rotatory + Con. 0.0001 61.68 1.04 42.31 0.97
-Beauty leaky Rotatory + Con. 0.0001 61.16 0.77 42.83 0.48
+Beauty silu Rotatory + Con 0.0001 61.68 1.04 42.31 0.97
+Beauty leaky Rotatory + Con 0.0001 61.16 0.77 42.83 0.48
 Men silu RMHA-4-Longer 0.0001 70.13 0.42 46.41 4.18
@@
 Fashion leaky RMHA-4-Longer 0.0001 76.20 0.51 49.85 3.92
-Games leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79
-Games silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55
-Games leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93
+Games leaky Rotatory + Con NaN 80.62 0.55 56.07 0.79
+Games silu Rotatory + Con-Longer NaN 80.29 0.47 55.18 0.55
+Games leaky Rotatory + Con-Longer NaN 80.21 1.51 55.06 1.93
 Games silu Learnt NaN 79.82 1.88 54.30 2.95
-Table 2: The table displays the top $4$ results for each dataset on cases with a standard deviation for $HIT$ below $3$.
-Dataset Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg
+Table 5. The table displays the top $4$ results for each dataset on cases with a standard deviation for Hit Meanbelow $3$.
+Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
 Beauty leaky Abs + Con 0.0001 67.93 4.17 48.71 3.59
@@
 Men leaky Learnt + Con 0.0001 71.34 7.13 57.92 13.07
-Men leaky Rotatory + Con. 0.1000 70.92 3.92 59.84 4.88
+Men leaky Rotatory + Con 0.1000 70.92 3.92 59.84 4.88
 Fashion silu RMHA-4 0.0001 77.26 0.24 49.75 2.33
@@
 Fashion leaky RMHA-4-Longer 0.0001 76.20 0.51 49.85 3.92
-Games leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79
-Games silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55
-Games leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93
+Games leaky Rotatory + Con NaN 80.62 0.55 56.07 0.79
+Games silu Rotatory + Con-Longer NaN 80.29 0.47 55.18 0.55
+Games leaky Rotatory + Con-Longer NaN 80.21 1.51 55.06 1.93
 Games silu Learnt NaN 79.82 1.88 54.30 2.95
-Table 3: he table displays the top $4$ results for each dataset on cases with a standard deviation for HIT below $12$.
-4 Experimental Setup
-4.1 Datasets
-We selected the Amazon dataset to assess the proficiency of various encoding types.
-For this purpose, we utilised four different real-world datasets extracted from product reviews on Amazon.com.
-These datasets have been diverse and widely used for SRS under the leave-one-out protocol (He and McAuley (2016b), Kang and McAuley (2018b), Hou et al.
-(2019), Wu et al.
-(2020), Steck (2019) and Zhou et al.
-(2020)).
-Since multiple versions of these datasets exist, we opted for the preprocessed versions from Rashed et al.
-(2022).
-These datasets encompass various product categories, providing a diverse and comprehensive source for training and evaluating recommendation models.
-They provide essential features for each interaction, such as user IDs, item IDs, timestamps, and additional vector contextual information.
-As we can see in Tables 5, and 4, they provide diversity in the sparsity, which is going to be a key point in our analysis, but also on the discreteness of the features.
-Dataset Users Items Interactions Attributes
-Men 34,244 110,636 254,870 2,048
-Fashion 45,184 166,270 358,003 2,048
-Games 31,013 23,715 287,107 506
-Beauty 52,204 57,289 394,908 6,507
-Table 4: Dataset Statistics.
-1.Beauty
-(Wu et al.
-(2020), Zhou et al.
-(2020) and Kang and McAuley (2018b)) The dataset includes discrete and categorical attributes for all beauty products, including fine-grained categories and brands.
-Mainly categorical and discrete features.
-2.Video Games
-(Wu et al.
-(2020) and Kang and McAuley (2018b)) The Video Games sub-dataset includes user interactions, reviews, and product details specific to the video game category as price, brand and categorical features.
-Most of the attributes here are discrete and categorical.
-3.Men
-(He and McAuley (2016b)) The men’s dataset encompasses a comprehensive collection of items falling under men’s clothing.
-The attributes are dense vectors from image-based features extracted from the last layer of a ResNet50 Deng et al.
-(2009) on the ImageNet dataset He et al.
-(2016a).
-3.Fashion
-(He and McAuley (2016b) and Hou et al. (2019)) It contains 6 categories for men’s and women’s clothing. The features were extracted as dense using the same ResNet50 approach.
-4.2 Models and Training
-To implement CARCA, we converted TensorFlow code from the original model into Pytorch [^1] https://github.com/researcher1741/Position_encoding_SRS.
-We extracted the implementation in Papso (2023) for the part of the loss.
-The original version in TensorFlow was written on top of previous models written in TensorFlow.
-We found that these implementations check the metrics every $20$ epoch, assuming that the overfitting starts later.
-This is probably because some models need more than $1000$ epochs, and the cost of measurement becomes significant.
-We measure the metrics in a logarithmic manner, giving rise to slightly better results in some cases.
-Our implementation can be found here .
-As in previous works (Rashed et al.
-(2022), Kang and McAuley (2018a), etc.), we use an ADAM optimiser to minimise the binary cross-entropy loss of the CARCA model while masking the padded items to prevent them from contributing to the loss function.
-For a given sequence of items interactions for a user $u$ as $\{i^{u}_{1},\ldots,i^{u}_{N}\}$, we create an input list as $I^{u+}=\{i^{u}_{1},\ldots,i^{u}_{N-1}\}$ by deleting the last item of the list, a positive target list as $T^{u+}=\{i^{u}_{2},\ldots,i^{u}_{N}\}$ by shifting one the input list, and a negative target list as $T^{u-}=\{i^{rand_{1}},\ldots,i^{rand_{N}-1}\}$ generated by random negative items.
-Then the loss is given by:
-$-\sum_{u\in U}\sum_{r\in T^{u+}\cup T^{u-}}\left[Y_{r}\log(\hat{Y}_{r})+(1-Y_{r})\log(1-\hat{Y}_{r})\right]$
-Here, $U$ represents the set of users, $T^{u+}\cup T^{u-}$ denotes the union of positive and negative items for user $u$, $Y_{r}$ is the observed interaction label for item $r$, and $\hat{Y}_{r}$ is the predicted probability of interaction with item $r$.
-4.3 Metrics
-We use two commonly used metrics for Next-Item Recommendations/SRS.
-Hit Rate measures the proportion of recommended items that overlap with the user’s actual interactions.
-Hit@10 counts the fraction of times the ground truth next item is among the top 10 items.
-The second metric is Normalized Discounted Cumulative Gain (NDCG). NDCG considers the relevance of items and their positions in the recommended list.
-$NDCG=\frac{DCG}{IDCG}$ (11)
-Where $DCG$ (Discounted Cumulative Gain) is calculated as the sum of the relevance scores of items at each position in the ranked list, discounted by their position, and $IDCG$ (Ideal Discounted Cumulative Gain) represents the maximum possible $DCG$ for a perfect ranking.
-$DCG=\sum_{i=1}^{n}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$ (12)
-$IDCG=\sum_{i=1}^{n}\frac{2^{max(rel_{i},0)}-1}{\log_{2}(i+1)}$ (13)
-Here, $rel_{i}$ is the item’s relevance at position $i$ in the ranking.
-5 Results
-We present all the results for all the datasets and encodings in Appendix A.1. Due to space limitations, sub-tables were extracted from them.
-Together with position encoding, we wanted to analyse other factors that could affect stability.
-In CARCA Rashed et al.
-(2022), they found $0.0001$ as the best upper bound ($nmax$ in our Tables) for the encoding vectors in all the datasets except for Games.
-We try different maximum values for all the encoding vectors ($0.1$ and $0.0001$) to see if their choice depends on the encoding.
-Except for some exceptions, the upper bound $0.0001$ produces better results.
-We also tried the activation function $silu$ instead of the $leakyrelu$.
-The results are slightly better regarding stability but without a strong correlation.
-5.1 Stability
-A common practice among the papers is to present an average of $5$.
-Our first experiments showed a significant difference between each one of the runs.
-We decided to analyse the deviation for the different trainings.
-The deviation and mean can be found in all the training tables.
-We also decided to compute the $95\%$ significance confidence interval and present it with the confidence interval in the complete tables results in Appendix A.1.
+Table 6. The table displays the top $4$ results for each dataset on cases with a standard deviation for Hit Meanbelow $12$.
+4.3. Metrics
+We employ two widely used metrics for Next-Item Recommendation and SRS: Hit@10 and Normalised Discounted Cumulative Gain (NDCG).
+Hit@10: This metric measures the fraction of times the ground truth next item appears within the top $10$ recommended items.
+For simplicity, we refer to Hit@10 as Hit, its mean as Hit Mean, its devisation as Hit Dev  and the average of the deviations as Avg Dev Hit.
+NDCG: Normalised Discounted Cumulative Gain evaluates the ranking quality by considering both the relevance of items and their positions in the recommended list. It is defined as:
+(14) $NDCG=\frac{DCG}{IDCG}$
+where $DCG$ (Discounted Cumulative Gain) is calculated as the sum of the relevance scores of items at each position in the ranked list, discounted by their position:
+(15) $DCG=\sum_{i=1}^{n}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$
+and $IDCG$ (Ideal Discounted Cumulative Gain) represents the maximum possible $DCG$ for a perfectly ranked list:
+(16) $IDCG=\sum_{i=1}^{n}\frac{2^{max(rel_{i},0)}-1}{\log_{2}(i+1)}$
+Here, $rel_{i}$ denotes the relevance score of the item at position $i$ in the ranking.
+We will denote its mean as NDCG Mean, its deviation as NDCG Dev  and the average of the deviations as Avg Dev NDCG.
+5. Results
+We present comprehensive results for all datasets and encodings in Appendix A.1.
+Due to space limitations, only selected sub-tables are included in this section.
+Furthermore, Appendix A.8 presents an additional experiment using SASRec++, confirming the consistency of our results.
+Details about the resources used for these experiments are provided in Appendix A.3.
+In addition to positional encoding, we analyze factors, such as maximum values and activation functions, that may influence model stability.
+In CARCA (Rashed et al., 2022), an upper bound of $0.0001$ ($nmax$ in our Tables) for the encoding vectors was identified as optimal across all datasets except for Games.
+To investigate whether the optimal upper bound varies with different encodings, we experimented with maximum values of $0.1$ and $0.0001$.
+Generally, an upper bound of $0.0001$ yielded better results, with a few exceptions.
+Furthermore, we compared activation functions by replacing $leakyrelu$ with $silu$.
+The $silu$ activation function resulted in slight improvements in stability; however, the correlation was not strong.
+5.1. Stability
+A common practice in the literature is to present the average results of $5$ runs with different random seeds.
+However, our preliminar experiments revealed significant variability between individual runs.
+To better understand this inconsistency, we analyzed the standard deviation across different trainings.
+Detailed deviation and mean values are provided in all the training tables.
+Additionally, we computed the $95\%$ confidence intervals for these metrics and included them in the complete results tables in Appendix A.1.
 Inter-Dataset Deviation:
-The stability varies from one dataset to another, being the Beauty dataset the most stable with an average Hit deviation of $0.9$ and NDCG deviation of $0.87$.
-Meanwhile, the maximum deviations are obtained by fashion.
-We can see that those datasets with a lower sparsity tend to be more stable, as do the Beauty and Games datasets, while those with a higher sparsity are less stable, as are the Men and Fashion datasets.
+The stability of results varies across different datasets.
+Beauty is the most stable, showing an average Hit deviation of $0.9$ and an average NDCG deviation of $0.87$, while Fashion has the highest deviations.
+Datasets with lower sparsity, such as Beauty and Games, generally exhibit greater stability.
+In contrast, datasets with higher sparsity, like Men and Fashion, show reduced stability.
+Additionally, there is a correlation with data type: lower-sparsity datasets often use discrete embedding representations.
+As discussed in Section 5.3, this stability pattern persists even when we adjust sparsity levels.
+By reducing datasets to achieve comparable sparsity, we show that both encoding selection and stability challenges remain for each data type.
 Inter-Encoding Deviation:
-In Table 6 and Table 12 in the Appendix, we see that some encodings produce more stable results than others.
-More precisely, the concatenated version for Abs , Learnt and Rotatory + Con.
-generally present more stable results.
-Furthermore, for the more stable case of $0.0001$ upper bound, in Table 6, Abs and Learnt are still significantly stable.
-At the same time, our Rotatory encoding produces similar results to the concatenated version.
-This could be because the average deviation is similar to no encoding, None .
-The most notable case is RMHA-4 , which, combined with the upper bound $0.0001$, gives the most stable results, around $0.51$, compared with the no-encoding case, $2.36$.
-This pattern is preserved to the general upper bound, suggesting that this encoding provides excellent stability.
-As we will see below, this encoding is worth trying in cases of high sparsity.
+Tables 4 and 19 (Appendix) reveal that some encodings yield more stable results than others.
+In particular, the concatenated versions of Abs , Learnt , and Rotatory demonstrate greater stability (Table 19).
+Under the most stable scenario with an upper bound of $0.0001$, as shown in Table 4, Abs + Con and Learnt + Con remain significantly more stable.
+However, Rotatory + Con shows no significant difference here, presenting deviations similar to None , likely due to the need to learn the angles.
+The most notable encoding for stability is RMHA-4 , which, combined with the $0.0001$ upper bound, produces the most stable results, with an Avg Dev Hit  of $0.51$ compared to $2.36$ for the no-encoding case.
+This pattern persists even with higher upper bounds, suggesting that RMHA-4 offers excellent stability, which can also be observed in the loss graph, as detailed in Appendix A.6.
+As we will discuss in Section 5.3, this encoding is particularly worth considering in unstable scenarios.
 Encoding Ablation:
-We can appreciate that Learnt and the sinusoidal Abs tend to have a similar deviation, higher than not using any encoding.
-To know if the stability is provided by the fact that RMHA-4 introduce positional information in each transformer or being a RPE head, we trained Rope .
-The results show that it is a mid between RMHA-4 on the one hand and Learnt and Abs on the other hand.
-It does not add or reduce stability for None .
-We wanted to know if being Rope in a middle point is due to being in all the transformer blocks while not being a APE encoding.
-To this goal, we added RopeOne , an ablation which applies Rope but only in the first block.
-The results are not far away from each other.
-We can see in Table 12 that RopeOne is even slightly better, while in Table 6, they are almost the same, in this case, slightly worse.
-Then, we decided to create a new encoding, Rotatory , which uses the rotatory properties but is not in the attention heads but a vector encoding.
-The results show the same stability as Rope and its ablation.
-The stability of the RMHA-4 encoding can also be appreciated in the losses. In Appendix A.5.
-Dataset Hit Dev. NDCG Dev. CI-length Density/Sparsity-1
-Men 3.9143 8.2742 5.6686 $56.727\times 10^{-}$ 5
-Fashion 4.87 9.7814 9.2257 $54.765\times 10^{-}$ 5
-Games 1.124 0.952 1.616 $43.904\times 10^{-}$ 4
-Beauty 0.935 0.87 1.22 $41.320\times 10^{-}$ 4
-Table 5: Deviations among the different datasets without Encoding. This pattern is generally preserved along the different encodings.
-encoding dev HIT dev NDCG runs CI-length
-Abs 5.43 8.25 9.38 6.71
-Abs + Con 3.52 5.56 7.12 5.35
-Learnt 4.79 7.64 7.88 7.05
-Learnt + Con 3.17 6.36 6.62 4.42
-None 2.35 5.26 8.00 3.29
-RMHA-4 0.51 1.08 6.00 0.88
-Rope 2.38 5.45 5.00 4.04
-Rotatory 2.46 5.01 8.38 3.48
-Rotatory + Con. 2.48 5.18 6.75 3.49
-RopeOne 2.82 6.54 4.29 5.46
-Rope-Longer 3.33 8.15 5.25 5.99
-Rotatory-Longer 1.05 1.04 4.67 2.08
-Rotatory + Con.-Longer 0.99 1.24 6.00 1.59
-RMHA-4-Longer 0.71 4.00 6.00 1.14
-Table 6: Deviations among the different encodings. Relative encoding in all the heads, RMHA-4 , presents more stability.
-5.2 Best Results
+To assess whether the stability is due to RMHA-4  ’s positional information at each transformer block or its RPE  -based nature, we tested RoPE , an APE .
+The results rank RoPE between RMHA-4 and Learnt /Abs , showing no significant change in stability compared to None (Table 4).
+We hypothesized that this lack of change in stability could be due to RoPE ’s application across all transformer blocks.
+To test this, we introduced RopeOne  , which applies RoPE only in the first block.
+Results in Table 19 show that RopeOne slightly outperforms RoPE , while Table 4 shows similar performance with a slight lower stability.
+This suggests that positional information at each transformer block does not significantly impact stability.
+Finally, we introduced Rotatory , which incorporates rotatory properties in a vector encoding.
+The results showed similar stability to RoPE and its ablation RopeOne  .
+This led us to the conclusion that the rotatory nature of some encodings and the relative encoding nature of RMHA-4 contribute to the stability of the experiments.
+5.2. Best Results
 Which is the best encoding?
-It depends on the spacing and the definition that one manages to get a good result.
-If we look at the table 7, we see that this dataset has excellent stability, but in cases like Learnt with $0.1$, with leakyrelu, the model gets over $74$ in HIT, but the $95\%$ confidence interval is ($60.74$, $88.24$).
-In other words, applying this approach cements the results of what was studied here Picard (2021); in some cases, it is better to get a good set of seeds than a suitable method.
+The optimal encoding depends on the maximum deviation considered.
+For instance, examining the Beauty dataset in Table 9, we observe lower deviations.
+However, in cases like Learnt with $nmax=0.1$ and leakyrelu, despite achieving $74$ in Hit Mean, the Hit Dev  increases to $17.19$.
+This highlights the trade-off between performance and stability, as some encodings may offer better performance metrics at the cost of higher instability.
+These findings support the conclusions of (Picard, 2021), which suggest that, in certain scenarios, selecting a robust set of random seeds can be more important than choosing the best encoding method.
 Maximum deviation of $12$:
-If we consider a maximum of $12$ as deviation, we get Table 3.
-In CARCA, they did an ablation study on the Men Dataset, showing that using Abs moves towards worse results.
-They produced similar results for Men, but both had a deviation of around $6$, so the results from the ablation were more a matter of initialisation.
-In this case, Men and Beauty get their best results with Learnt -like and Abs -like encodings, respectively.
-But they are unstable.
-In contrast, Fashion and Games get their top values with RMHA-4 and Rotatory , respectively, which are stable.
+When setting the maximum Hit Dev  to $12$, we get Table 6.
+In the CARCA (Rashed et al., 2022), an ablation analysis of the Men dataset revealed that using Abs results in poorer performance than None .
+Similar results were observed for the Men in our experiments, where both cases showed a deviation of around $6$, suggesting that the observed effects were more due to initialization than encoding choice.
+With a maximum of $12$ on Hit Dev, Men and Beauty achieve their best performance with Learnt -like and Abs -like encodings, respectively, although both are unstable.
+In contrast, Fashion and Games reach their optimal results with RMHA-4 and Rotatory , respectively, both of which are stable.
+Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
+Submen leaky RMHA-4 0.1000 77.28 2.66 68.45 3.65
+Submen leaky RMHA-4 0.0001 75.96 2.71 66.03 4.00
+Subfashion leaky RMHA-4 0.0001 79.72 2.85 71.57 3.66
+Subfashion leaky None 0.1000 32.94 1.89 16.77 3.00
+Subgames leaky Rotatory + Con NaN 52.23 1.57 31.51 0.96
+Subgames silu Rotatory + Con NaN 50.73 0.89 30.48 0.72
+Subgames leaky RMHA-4 NaN 50.43 0.95 29.96 0.87
+Subgames leaky None NaN 49.53 1.01 28.84 0.69
+Submen3 leaky Rotatory + Con 0.1000 73.12 2.38 62.86 2.99
+Submen3 silu Rotatory + Con 0.0001 72.77 2.96 62.56 4.31
+Table 7.
+The table displays the top $4$ results for each ablation dataset on cases with a Hit Devbelow $3$.
+In the case of the Fashion dataset ablation we did not get any training below $3$ in the deviation.
 Maximum deviation of $3$:
-If we reduce the maximum deviation to $3$, we get some results that we consider the right results.
-As we can see in Table 2, the best encodings are RMHA-4 and Rotatory (or its concatenated version), and they correlate with the sparsity.
-Those with a high sparsity benefit more from the extra stability provided by RMHA-4 , while those with a lower sparsity, Beauty and Games, get their best through a purely rotatory approach Rotatory .
-We can see that even considering only those which are stable, the results obtained in all the datasets are new SOTA.
-Encoding "Ablation":
-We created some ablations from existing encoding techniques to analyse where the stability came from.
-So far, our experiments used almost the same hyperparameters as the original CARCA.
-Looking again at the losses, we found that some of the top results did not converge, i.e., the loss and the metrics were improving.
-Thus, we decided to extend some trainings with more epochs to see if this produced better results.
-We added $400$ epochs to the default values (11).
-The improvements were not significant in this regard.
-6 Conclusion
-Our main conclusion is that encodings should not be considered minor in this model family.
-On the contrary, they are a powerful tool, and depending on the data set, a couple of possible better choices can be pre-selected.
-We have shown that the higher the sparsity, the better the results, as they stabilise the training until the best results are achieved.
-On the other hand, when the set density is sufficient, our rotatory encodings are a good choice or a learnable encoding.
-Also, at publication time, we have achieved SOTA results using only the encodings.
-And these results are, in most cases, more reliable than previous SOTA results.
-7 Future work
-During the encoding analysis, we found some correlation between the bound used for the encoding and the stability.
-Previous works have experimented with very low bounds, usually below $0.001$.
-We would like to go deep in this direction and analyse the behaviour of the metric used for encoding and its bounds.
-8 Limitations
-Despite the details of this study, it was carried out using a transformer architecture.
-Today, the latest models in the field are based on this architecture or GCNs.
-Using other architectures may affect this study and the conclusions on using encodings.
-The fact that the choice depends on the dataset can give a first idea of the possible encodings in case of using different architectures.
+Reducing the maximum Hit Dev  to $3$ yields optimal results, as presented in Table 5.
+In this Table, the best encodings are RMHA-4 and Rotatory (or its concatenated version), which correlate strongly with sparsity, datatype, and stability (low-deviation).
+As discussed in Section 5.3, datasets with high-deviation without encoding benefit from the stability provided by RMHA-4 , while low-deviation datasets like Beauty and Video Games perform best with the purely rotatory approach of Rotatory .
+Even among stable datasets, these results surpass the previous SOTA benchmarks (CARCA) at the time.
+We also observe a significant improvement over the None case for high-deviation datasets, with gains exceeding $5\%$, while achieving greater stability.
+Impact of Extended Training on Encoding Stability:
+Ablation studies were conducted on the encoding techniques to investigate the source of stability.
+Our initial experiments used hyperparameters similar to those of the original CARCA model (Appendix A.4).
+Upon reviewing the loss metrics, we found that several top-performing models did not converge, as both loss and performance metrics continued to improve.
+To address this, we extended the training by an additional $400$ epochs beyond the default values specified in Table 18.
+Despite this extension, the improvements in model performance were minimal, indicating that the optimizer may have been oscillating around a local minimum.
+5.3. Sparsity and Deviation Reduction
+As seen in Section 5.2, two encodings, RMHA-4 and Rotatory + Con , consistently provide the best results (with maximum Hit Dev  of $3$).
+There appears to be a correlation between the deviation of datasets and the choice of optimal encoding.
+For high-deviation datasets like Men and Fashion, RMHA-4 yields better results, whereas for low-deviation datasets like Games and Beauty, Rotatory + Con performs best.
+Importantly, both Men and Fashion share two other key characteristics: dense vector representations and higher-sparsity.
+This makes it difficult to attribute the performance solely to the default dataset deviation.
+The complete tables for the subset results are available in Appendix A.2.
+Sparsity Reduction:
+To investigate this further, We conducted a sparsity reduction experiment by creating subdatasets with densities comparable to other datasets.
+Specifically, we reduced the sparsity of Fashion from $4.765\cdot 10^{-5}$ to $3.564\cdot 10^{-4}$ and Men from $6.727\cdot 10^{-5}$ to $2.860\cdot 10^{-4}$, resulting in Subfashion and Submen.
+We applied RMHA-4 , Rotatory + Con , and None to these datasets.
+Table 7 shows that RMHA-4 consistently produced the best results across both subdatasets, confirming that its effectiveness is not solely due to sparsity.
+In Subfashion, Hit Devincreased to $15.105$, and NDCG Devto $17.935$, while Submen experienced a small reduction in deviations (Hit Dev7.2967.2967.2967.296, NDCG Dev9.0969.0969.0969.096).
+Although Men saw a slight reduction on deviation (Table 8), it wasn’t substantial.
+Additionally, we created Subgames for comparison, reducing the Games dataset’s sparsity to see if it influences performance.
+Interestingly, Rotatory + Con remained the best encoding for this more stable dataset (Table 7), suggesting that even after sparsity reduction, dataset characteristics like deviation may drive encoding performance.
+Deviation Reduction:
+To rule out the correlation with the vector representation and isolate the effect of deviation, we further reduced the Men dataset to $10k$ users and $80k$ items ($10000$ users and $45129$ items after filtering).
+This reduced sparsity to $1.648\cdot 10^{-4}$ and helped stabilize the experiments, decreasing Hit Devto $5.198$ and NDCG Devto $6.826$.
+The reduced deviation was enough to make Rotatory + Con the best-performing encoding.
+Although the performance is close to that of the None case (Table 17), there was a shift in the optimal encoding.
+These findings indicate that high-deviation datasets, even after sparsity reduction, benefit from the added stability of RMHA-4 .
+In contrast, low-deviation datasets like Subgames, perform better with Rotatory + Con  .
+Dataset Users Items Interac. Hit Dev NDCG Dev CI-length Density
+Subfashion 3,840 5,000 16,080 15.105 17.935 24.175 $3.564\cdot 10^{-4}$
+Submen 8,332 10,000 28,603 7.2925 9.0925 10.515 $2.860\cdot 10^{-4}$
+Subgames 5,000 4,601 33,353 1.115 0.775 1.79 $1.3341\cdot 10^{-3}$
+Submen 2 10,000 45,129 74,391 5.1975 6.8225 8.31 $1.648\cdot 10^{-4}$
+Table 8. Statistics of the ablations: Deviations for Hit@10 and NDCG without Encoding. Together with the average length of the $CI$ intervals, and the density.
+6. Conclusion
+In this study, we emphasize that encodings are not mere technical details but rather critical components that can significantly influence both the performance and stability of SRS.
+Our findings reveal that high-deviation datasets benefit from relative encodings, such as RMHA-4 , which not only stabilize training but also yield optimal results.
+These encodings play a vital role in mitigating performance fluctuations, allowing the model to consistently perform well across different runs.
+On the other hand, for low-deviation datasets, our proposed Rotatory encoding offers a robust alternative that further improves results without introducing unnecessary complexity.
+Importantly, at the time of publication, we achieved state-of-the-art performance using only these encoding methods, surpassing previous benchmarks and demonstrating greater reliability.
+Our results indicate that encodings are a pivotal factor in ensuring the stability of transformer-based models, especially in recommendation systems.
+We encourage the research community to place greater emphasis on positional encodings, as they can drastically enhance model effectiveness, particularly in tasks involving sequential data.
+In the Appendix A.7, there is a decision tree for the encoding election.
 Appendix A Appendix
-A.1 Experiments
+A.1. Experiments
 In this section, we present all our results for the four datasets, thereby completing the missing experiments from the sub-tables in the main corpus.
-In tables 7, 8, and 9, we can find Rope-Longer , which is not presented in Table 1.
+In tables 9, 10, and 11, we can find RoPE-Longer , which is not presented in Table 1.
 These are just a few cases in which we extend the number of epochs since we found that for RoPE , it takes longer to reach its maximum.
-Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
+Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
 leaky None 0.0001 56.00 0.96 37.93 0.65 9 (55.37, 56.63) 1.26
@@
 silu Abs + Con 0.1000 54.54 3.33 35.43 3.34 5 (51.62, 57.46) 5.84
-leaky Rotatory + Con. 0.0001 61.16 0.77 42.83 0.48 3 (60.29, 62.03) 1.74
-silu Rotatory + Con. 0.0001 61.68 1.04 42.31 0.97 3 (60.5, 62.86) 2.36
-leaky Rotatory + Con. 0.1000 50.51 1.59 32.25 1.76 3 (48.71, 52.31) 3.60
-silu Rotatory + Con. 0.1000 50.88 1.00 32.70 0.87 3 (49.75, 52.01) 2.26
+leaky Rotatory + Con 0.0001 61.16 0.77 42.83 0.48 3 (60.29, 62.03) 1.74
+silu Rotatory + Con 0.0001 61.68 1.04 42.31 0.97 3 (60.5, 62.86) 2.36
+leaky Rotatory + Con 0.1000 50.51 1.59 32.25 1.76 3 (48.71, 52.31) 3.60
+silu Rotatory + Con 0.1000 50.88 1.00 32.70 0.87 3 (49.75, 52.01) 2.26
 leaky Learnt 0.0001 59.83 3.71 39.85 3.11 3 (55.63, 64.03) 8.40
@@
 silu Rotatory 0.1000 49.40 0.96 31.14 1.04 5 (48.56, 50.24) 1.68
-Table 7: Beauty results with different positional encodings. Original results from Rashed et al. (2022) at the end.
-Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
+Table 9. Beauty results with different positional encodings. Original results from (Rashed et al., 2022) at the end.
+Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
 leaky None 0.0001 65.44 3.54 46.46 9.99 9 (63.13, 67.75) 4.62
@@
 silu Abs + Con 0.1000 65.14 8.20 51.71 11.06 13 (60.68, 69.6) 8.92
-leaky Rotatory + Con. 0.0001 69.13 2.80 55.33 7.69 9 (67.3, 70.96) 3.66
-silu Rotatory + Con. 0.0001 71.16 1.38 48.61 8.62 7 (70.14, 72.18) 2.04
-leaky Rotatory + Con. 0.1000 68.61 2.76 57.08 3.72 7 (66.57, 70.65) 4.08
-silu Rotatory + Con. 0.1000 65.23 7.76 52.99 9.31 11 (60.64, 69.82) 9.18
+leaky Rotatory + Con 0.0001 69.13 2.80 55.33 7.69 9 (67.3, 70.96) 3.66
+silu Rotatory + Con 0.0001 71.16 1.38 48.61 8.62 7 (70.14, 72.18) 2.04
+leaky Rotatory + Con 0.1000 68.61 2.76 57.08 3.72 7 (66.57, 70.65) 4.08
+silu Rotatory + Con 0.1000 65.23 7.76 52.99 9.31 11 (60.64, 69.82) 9.18
 leaky Learnt 0.0001 68.37 5.75 51.41 12.43 13 (65.24, 71.5) 6.26
@@
 leaky None 0.0001 59.1 38.1 5
-Table 8: Fashion Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
-Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
+Table 10. Fashion Dataset results with different positional encodings. Results from (Rashed et al., 2022) at the end.
+Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
 leaky None 0.0001 65.19 6.37 48.99 11.96 9 (61.03, 69.35) 8.32
@@
 silu Abs + Con 0.1000 63.33 7.50 48.62 10.54 9 (58.43, 68.23) 9.80
-leaky Rotatory + Con. 0.0001 70.64 5.29 59.30 6.98 10 (67.36, 73.92) 6.56
-silu Rotatory + Con. 0.0001 67.18 5.82 50.27 13.21 10 (63.57, 70.79) 7.22
-leaky Rotatory + Con. 0.1000 70.92 3.92 59.84 4.88 6 (67.78, 74.06) 6.28
-silu Rotatory + Con. 0.1000 63.88 7.42 50.69 9.75 10 (59.28, 68.48) 9.20
+leaky Rotatory + Con 0.0001 70.64 5.29 59.30 6.98 10 (67.36, 73.92) 6.56
+silu Rotatory + Con 0.0001 67.18 5.82 50.27 13.21 10 (63.57, 70.79) 7.22
+leaky Rotatory + Con 0.1000 70.92 3.92 59.84 4.88 6 (67.78, 74.06) 6.28
+silu Rotatory + Con 0.1000 63.88 7.42 50.69 9.75 10 (59.28, 68.48) 9.20
 leaky Learnt 0.0001 72.08 4.44 56.55 9.98 7 (68.79, 75.37) 6.58
@@
 leaky None 0.0001 55.0 34.9 5
-Table 9: Men Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
-Act encoding nmax mean_hit dev_hit mean_ndcg dev_ndcg runs CI CI-length
+Table 11. Men Dataset results with different positional encodings. Results from (Rashed et al., 2022) at the end.
+Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
 leaky None NaN 78.95 1.25 53.37 1.21 6 (77.95, 79.95) 2.00
@@
 silu Learnt + Con NaN 79.32 1.50 53.68 1.97 6 (78.12, 80.52) 2.40
-leaky Rotatory + Con. NaN 80.62 0.55 56.07 0.79 6 (80.18, 81.06) 0.88
-silu Rotatory + Con. NaN 79.61 2.17 54.80 2.71 6 (77.87, 81.35) 3.48
+leaky Rotatory + Con NaN 80.62 0.55 56.07 0.79 6 (80.18, 81.06) 0.88
+silu Rotatory + Con NaN 79.61 2.17 54.80 2.71 6 (77.87, 81.35) 3.48
 leaky Abs + Con NaN 79.71 1.28 54.93 1.50 6 (78.69, 80.73) 2.04
@@
 silu Learnt NaN 79.82 1.88 54.30 2.95 8 (78.52, 81.12) 2.60
-leaky Rotatory + Con.-Longer NaN 80.21 1.51 55.06 1.93 6 (79.0, 81.42) 2.42
-silu Rotatory + Con.-Longer NaN 80.29 0.47 55.18 0.55 6 (79.91, 80.67) 0.76
+leaky Rotatory + Con-Longer NaN 80.21 1.51 55.06 1.93 6 (79.0, 81.42) 2.42
+silu Rotatory + Con-Longer NaN 80.29 0.47 55.18 0.55 6 (79.91, 80.67) 0.76
 leaky Rotatory-Longer NaN 78.00 1.10 52.30 1.05 6 (77.12, 78.88) 1.76
@@
 silu Rotatory NaN 76.21 1.77 51.06 2.04 6 (74.79, 77.63) 2.84
-leaky None 0.0001 78.2 57.3 5
-Table 10: Games Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
-A.2 Resources
-Due to the instability of these models, we had to run many seeds to get some confident results.
-We started by running each one of the model configurations three times.
-After that, we decided to run more seeds based on the confidence of each one of the cases.
-We tried to keep at least a confidence range of ten points.
-We used $V100$ GPUs of 32GBs.
-For $men$ and $fashion$, based on the encoding, the training times go from $4$ to $6$ hours.
-For $Beauty$, they went from $22$ to $26$ hours.
-Based on the team’s resource ability, these experiments were carried out over $6$ months, lasting between $1$ and $8$.
-We used the following seeds: $42$, $43$, $44$, $45$, $46$, $1809$, $1810$, $1811$, $1812$, $1741$, $1742$, $1743$, $1744$, $1745$, $123$, $234$, $345$, $456$, $567$, $678$, $789$, $890,102$, $203$, $304$, $405$, $506$, $607$, $708$, $808$, $901$, $1001$, $987$, $8765$, $7654$, and $6543$.
-A.3 Hyperparameters
-In this section, we present the hyperparameters used for training.
-As mentioned above, we use the same hyperparameters as CARCA Rashed et al.
-(2022).
-In some cases, we experimented with the upper bound from the norm in the encoding and the training length.
+leaky None NaN 78.2 57.3 5
+Table 12. Games Dataset results with different positional encodings. Results from (Rashed et al., 2022) at the end.
+A.2. Sparsity and deviation experiments
+In this section, we present tables of the datasets created by reduction to evaluate whether the encoding type depends on the vector representation of the items, deviation, and sparsity.
+For these cases, we considered only three encodings: RMHA-4 , Rotatory , and None .
+Table 17 contains results for datasets with a deviation below $12$, which we deemed invalid and have therefore moved to this appendix.
+Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
+leaky None 0.0001 48.00 22.19 33.53 26.31 6 (30.24, 65.76) 35.52
+silu None 0.0001 66.30 18.61 55.17 21.99 6 (51.41, 81.19) 29.78
+leaky None 0.1000 32.94 1.89 16.77 3.00 6 (31.43, 34.45) 3.02
+silu None 0.1000 63.85 17.73 52.02 20.44 6 (49.66, 78.04) 28.38
+leaky Rotatory + Con 0.0001 50.04 19.62 35.82 22.95 6 (34.34, 65.74) 31.40
+silu Rotatory + Con 0.0001 63.87 19.99 52.59 22.77 6 (47.87, 79.87) 32.00
+leaky Rotatory + Con 0.1000 49.17 21.02 35.89 25.10 6 (32.35, 65.99) 33.64
+silu Rotatory + Con 0.1000 67.94 19.26 57.24 23.51 6 (52.53, 83.35) 30.82
+leaky RMHA-4 0.0001 79.72 2.85 71.57 3.66 6 (77.44, 82.0) 4.56
+silu RMHA-4 0.0001 46.55 14.12 31.68 16.59 6 (35.25, 57.85) 22.60
+leaky RMHA-4 0.1000 57.99 20.07 45.36 24.01 6 (41.93, 74.05) 32.12
+silu RMHA-4 0.1000 61.14 14.64 48.46 18.10 6 (49.43, 72.85) 23.42
+Table 13. SubFashion Dataset results for the best positional encodings and the None encoding case.
+Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
+leaky None 0.0001 77.10 3.72 67.89 5.25 12 (75.0, 79.2) 4.20
+silu None 0.0001 76.33 6.11 67.00 7.08 12 (72.87, 79.79) 6.92
+leaky None 0.1000 69.29 15.21 58.64 18.19 6 (57.12, 81.46) 24.34
+silu None 0.1000 74.66 4.13 64.38 5.85 6 (71.36, 77.96) 6.60
+leaky Rotatory + Con 0.0001 71.82 11.90 61.69 14.11 12 (65.09, 78.55) 13.46
+silu Rotatory + Con 0.0001 75.28 3.63 65.31 5.03 12 (73.23, 77.33) 4.10
+leaky Rotatory + Con 0.1000 73.93 5.87 63.77 8.00 6 (69.23, 78.63) 9.40
+silu Rotatory + Con 0.1000 73.11 5.43 62.64 6.80 6 (68.77, 77.45) 8.68
+leaky RMHA-4 0.0001 75.96 2.71 66.03 4.00 8 (74.08, 77.84) 3.76
+silu RMHA-4 0.0001 70.44 9.61 59.67 11.50 8 (63.78, 77.1) 13.32
+leaky RMHA-4 0.1000 77.28 2.66 68.45 3.65 6 (75.15, 79.41) 4.26
+silu RMHA-4 0.1000 72.54 4.57 61.75 5.96 6 (68.88, 76.2) 7.32
+Table 14. SubMen Dataset results for the best positional encodings and the None encoding case.
+Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
+leaky None NaN 49.53 1.01 28.84 0.69 6 (48.72, 50.34) 1.62
+silu None NaN 48.31 1.22 28.36 0.86 6 (47.33, 49.29) 1.96
+leaky Rotatory + Con NaN 52.23 1.57 31.51 0.96 6 (50.97, 53.49) 2.52
+silu Rotatory + Con NaN 50.73 0.89 30.48 0.72 6 (50.02, 51.44) 1.42
+leaky RMHA-4 NaN 50.43 0.95 29.96 0.87 6 (49.67, 51.19) 1.52
+silu RMHA-4 NaN 48.97 1.35 28.57 0.79 6 (47.89, 50.05) 2.16
+Table 15. Subgames Dataset results for the best positional encodings (Rotatory + Con and RMHA-4 ) and the None encoding case.
+Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
+leaky None 0.0001 72.08 6.04 61.47 7.96 6 (67.25, 76.91) 9.66
+silu None 0.0001 73.95 3.33 64.05 4.71 6 (71.29, 76.61) 5.32
+leaky None 0.1000 71.56 5.38 60.96 6.93 6 (67.26, 75.86) 8.60
+silu None 0.1000 70.51 6.04 59.29 7.69 6 (65.68, 75.34) 9.66
+leaky Rotatory + Con 0.0001 71.52 6.39 60.71 8.17 6 (66.41, 76.63) 10.22
+silu Rotatory + Con 0.0001 72.77 2.96 62.56 4.31 6 (70.4, 75.14) 4.74
+leaky Rotatory + Con 0.1000 73.12 2.38 62.86 2.99 6 (71.22, 75.02) 3.80
+silu Rotatory + Con 0.1000 72.03 3.90 61.33 4.76 6 (68.91, 75.15) 6.24
+silu RMHA-4 0.0001 53.67 8.81 36.15 12.73 12 (48.69, 58.65) 9.96
+leaky RMHA-4 0.1000 61.82 12.74 48.93 14.82 6 (51.63, 72.01) 20.38
+silu RMHA-4 0.1000 61.27 5.87 47.24 6.96 6 (56.57, 65.97) 9.40
+Table 16. SubMen2 Dataset results for the best positional encodings and the None encoding case.
+Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
+Submen leaky RMHA-4 0.1000 77.28 2.66 68.45 3.65
+Submen leaky None 0.0001 77.10 3.72 67.89 5.25
+Submen silu None 0.0001 76.33 6.11 67.00 7.08
+Submen leaky RMHA-4 0.0001 75.96 2.71 66.03 4.00
+Subfashion leaky RMHA-4 0.0001 79.72 2.85 71.57 3.66
+Subfashion leaky None 0.1000 32.94 1.89 16.77 3.00
+Subgames leaky Rotatory + Con NaN 52.23 1.57 31.51 0.96
+Subgames silu Rotatory + Con NaN 50.73 0.89 30.48 0.72
+Subgames leaky RMHA-4 NaN 50.43 0.95 29.96 0.87
+Subgames leaky None NaN 49.53 1.01 28.84 0.69
+Submen3 silu None 0.0001 73.95 3.33 64.05 4.71
+Submen3 leaky Rotatory + Con 0.1000 73.12 2.38 62.86 2.99
+Submen3 silu Rotatory + Con 0.0001 72.77 2.96 62.56 4.31
+Submen3 leaky None 0.0001 72.08 6.04 61.47 7.96
+Table 17. The table displays the top $4$ results for each dataset on cases with a standard deviation for HIT below $12$.
+A.3. Resources
+Due to the inherent instability of these models, it was necessary to run multiple seeds to obtain reliable results.
+Initially, we executed each model configuration three times.
+However, as we analyzed the results, we determined that additional seeds were needed for cases with high-deviation, aiming to maintain a confidence interval within ten points.
+All experiments were conducted on $V100$ GPUs with $32$ GB of memory.
+Training times varied depending on the dataset and encoding type.
+For the Men and Fashion datasets, training durations ranged from $4$ to $6$ hours, while for Beauty, the training times extended from $22$ to $26$ hours.
+Given the available resources, these experiments were spread over a period of six months.
+We used the following seed values to ensure diversity in our training runs: $42$, $43$, $44$, $45$, $46$, $1809$, $1810$, $1811$, $1812$, $1741$, $1742$, $1743$, $1744$, $1745$, $123$, $234$, $345$, $456$, $567$, $678$, $789$, $890,102$, $203$, $304$, $405$, $506$, $607$, $708$, $808$, $901$, $1001$, $987$, $8765$, $7654$, and $6543$.
+A.4. Hyperparameters
+In this section, we outline the hyperparameters used during training.
+Consistent with the approach used in CARCA (Rashed et al., 2022), we adopted the same baseline hyperparameters across most experiments to maintain comparability.
+In specific cases, we experimented with variations in the upper bound of the norm for the encodings, activation functions, as well as adjusting the training duration to explore its impact on stability and performance.
 Hyperparameter Men Fashion Games Beauty
@@
 Residual connection in the cross-attention block No No Yes Yes
-Table 11: Hyperparameters with corresponding options were used in the original CARCA model.
+Table 18.
+Hyperparameters with corresponding options were used in the original CARCA model.
 This parameter were written in tensor flow.
@@
 weight", the value 00 corresponds to no bound, which in Pytorch is $Nan$, and we denote it as None .
-A.4 Deviations over all the trainings
-Here, we present the table with all the deviation values from the training.
-In the table in the main text, we considered those training where we use the default values for the upper bound, i.e., $0.0001$, and $None$ for the Games dataset.
-We can see that the trends pointed out in Section 5.1 are still valid in this table.
-The encodings with concatenation are more stable, and RMHA-4 is the most stable one.
-We can see that with longer trainings Rotatory encodings become more stable than Rope
-encoding HIT dev NDCG dev runs CI-length
+A.5. Deviations over all the trainings
+In this section, we present the complete table of deviation values from all training sessions.
+In the main text, we focused on the training runs where the default upper bound values ($0.0001$) were used, along with $None$ for the Games dataset.
+The trends discussed in Section 5.1 remain consistent here.
+Encodings that incorporate concatenation generally demonstrate greater stability, with RMHA-4 continuing to be the most stable overall.
+Additionally, we observe that with extended training durations, the Rotatory encodings surpass RoPE in stability.
+encoding Avg Dev Hit Avg Dev NDCG runs CI-length
 Abs 7.20 9.42 10.57 8.42
@@
 Rotatory 3.92 6.10 9.00 4.92
-Rotatory + Con. 3.16 5.12 6.71 4.47
+Rotatory + Con 3.16 5.12 6.71 4.47
 RopeOne 2.82 6.54 4.29 5.46
@@
 Rotatory-Longer 1.31 1.35 4.00 3.00
-Rotatory + Con.-Longer 0.99 1.24 6.00 1.59
+Rotatory + Con-Longer 0.99 1.24 6.00 1.59
 RMHA-4-Longer 0.71 4.00 6.00 1.14
-Table 12: Deviations among the different encodings. Those per head, like RMHA-4 or Rope , are more stable than others.
-A.5 Losses
+Table 19. Deviations among the different encodings. Those per head, like RMHA-4 or RoPE , are more stable than others, together with Rotatory .
+A.6. Losses
 This section shows some loss cases from our trainings showing the stability of RMHA-4 .
@@
 The second image shows the validation Losses for the analogous case in the Men Dataset.
-Figure 2: Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ ad silu. Dataset: Fashion
-Figure 3: Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ ad silu. Dataset: Men
+Figure 4. Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ and silu. Dataset: Fashion
+Figure 5. Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ and silu. Dataset: Men
+A.7. Decision diagram
+Figure 6. Decision tree for encoding selection based on dataset deviation. The diagram outlines the recommended encodings.
+A.8. Extra experiments with SASRec++
+All of our experiments were conducted using the CARCA model (Rashed et al., 2022).
+To assess whether the encoding behaviors observed in CARCA extend to other architectures, we also evaluated the SASRec++ model, an extension of the original SASRec model (Kang and McAuley, 2018a).
+SASRec+++++ + is conceptually similar to CARCA but differs in that it replaces the transformer decoder with a simple dot product operation.
+We applied SASRec++ to the Submen dataset using the following encodings: None , Learnt + Con , Rotatory + Con , Abs + Con , Learnt , Abs , RMHA-4 , RopeOne , RoPE , and Rotatory .
+As shown in Table 20, we find that RMHA-4 continues to outperform other encodings, maintaining its position as the best-performing encoding across all experiments.
+Most of the results cluster around $40\%$, with minimal differences observed between encodings.
+The difference between RopeOne and other encodings is particularly small, with only slight improvements seen for RMHA-4 with a low Hit Dev  and Learnt + Con (as seen in Table 21) with a higher Hit Dev.
+Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
+Submen silu RMHA-4 0.1000 44.12 2.58 19.73 1.17
+Submen silu RMHA-4 0.0001 42.45 2.91 18.91 1.11
+Submen leaky RopeOne 0.0001 40.69 0.75 18.90 0.60
+Submen leaky RopeOne 0.1000 40.66 0.74 18.67 0.39
+Table 20. The table displays the top $4$ results for Submen for SASRec++ on cases with a standard deviation for Hit Meanbelow $3$.
+Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
+Submen silu Learnt + Con 0.0001 58.08 11.49 26.41 6.37
+Submen silu RMHA-4 0.1000 44.12 2.58 19.73 1.17
+Submen leaky RMHA-4 0.0001 42.86 5.52 20.57 5.34
+Submen silu RMHA-4 0.0001 42.45 2.91 18.91 1.11
+Table 21. The table displays the top $4$ results for Submen for SASRec++ on cases with a standard deviation for Hit Meanbelow $12$.